### v1

# Unveiling Downstream Performance Scaling of LLMs: A Clustering-Based Perspective

Abstract
The rapid advancements in computing dramatically increase the scale and cost of training Large Language Models (LLMs). Accurately predicting downstream task performance prior to model training is crucial for efficient resource allocation, yet remains challenging due to two primary constraints: (1) the “emergence phenomenon”, wherein downstream performance metrics become meaningful only after extensive training, which limits the ability to use smaller models for prediction; (2) Uneven task difficulty distributions and the absence of consistent scaling laws, resulting in substantial metric variability. Existing performance prediction methods suffer from limited accuracy and reliability, thereby impeding the assessment of potential LLM capabilities. To address these challenges, we propose a Clustering-On-Difficulty (COD) downstream performance prediction framework. COD first constructs a predictable support subset by clustering tasks based on difficulty features, strategically excluding non-emergent and non-scalable clusters. The scores on the selected subset serve as effective intermediate predictors of downstream performance on the full evaluation set. With theoretical support, we derive a mapping function that transforms performance metrics from the predictable subset to the full evaluation set, thereby ensuring accurate extrapolation of LLM downstream performance. The proposed method has been applied to predict performance scaling for a 70B LLM, providing actionable insights for training resource allocation and assisting in monitoring the training process. Notably, COD achieves remarkable predictive accuracy on the 70B LLM by leveraging an ensemble of small models, demonstrating an absolute mean deviation of 1.36% across eight important LLM evaluation benchmarks.
1 Introduction
Large Language Models (LLMs) have emerged as transformative technologies in natural language understanding, generation, and reasoning [1, 14, 5]. Their impressive success heavily relies on scaling model parameters and pre-training data, with training loss empirically following a power-law relationship with compute [18, 21]. However, this reduction in training loss primarily reflects an in-domain compression effect and does not necessarily indicate improved out-of-domain generalization or downstream performance–the factor of primary concern in practice. Specifically, performance scaling of downstream tasks aims to predict the accuracy of the target LLM on downstream tasks using metrics from smaller models. Our objective is to develop a prediction method that works reliably across a diverse range of downstream tasks, minimizing the worst-case prediction error.
Despite extensive efforts, a reliable scaling law for downstream tasks remains elusive. One line of work attempts to extrapolate large-model performance by modeling the performance-loss relationship [6, 13, 8, 38, 26], but this often fails to capture the emergent behaviors of LLMs and the mismatch between in-domain loss and downstream metrics [42]. Another line of research focuses on direct extrapolation of performance-compute relationship [1, 19], yet the uneven difficulty distribution across different evaluation samples undermines its accuracy. We observe that different evaluation samples actually follow distinct performance scaling patterns, and thus applying a single extrapolation formula to the entire evaluation set is suboptimal. We give the detailed analysis in Section 3.
To address these challenges, we propose a new performance scaling law, derived from the existing loss scaling law [21], specifically applicable to evaluation subsets that exhibit consistent performance scaling patterns. Building on the performance scaling law, we develop a Clustering-On-Difficulty (COD) multi-stage framework for predicting downstream performance. Specifically, we first cluster tasks by their difficulty features, and then filter out clusters that lack valid extrapolation patterns. Next, we fit the performance-compute relationships in the remaining clusters under our performance scaling law, extrapolate the performance of large models within these clusters, and finally map the aggregated predictions to the complete task set.
We validate our COD approach on eight evaluation sets, including popular MATH [15], BBH [32], and MMLU pro [36]. COD achieves an average prediction error of 1.36% on a 70B-parameter LLM. Our results demonstrate that this difficulty-aware framework substantially outperforms existing methods, establishing a promising paradigm for accurate downstream performance scaling of LLMs.
Our contributions can be summarized as follows:
  • We propose the COD framework to address high variance and emergent phenomena in the LLM performance by effectively modeling the difficulty distribution within the evaluation sets.
  • We introduce a performance scaling law for cluster-wise performance prediction, with theoretical support and experimental validation.
  • Extensive experiments conducted across eight diverse evaluation sets demonstrate that COD achieves a state-of-the-art average prediction error of 1.36% on a 70B-parameter LLM.
2 Related Work
2.1 Loss Scaling Laws
Loss scaling laws provide a systematic framework for understanding the relationship between computational resources, data, model size, and the final performance of LLMs. Early work by Kaplan et al. [21] demonstrates that the pre-training loss of LLMs follows a power-law relationship with the compute (the number of floating-point operations) used in training. Subsequent studies extend these findings to other domains, such as computer vision [41], graph learning [24] and vision-language models [2, 16]. Recent research has also explored scaling laws in specific contexts, such as fine-tuning [17, 34], vocabulary size optimization [33], retrieval-augmented models [30], and hyperparameter tuning [23, 40]. These studies highlight the broad applicability of scaling laws and their potential to guide the efficient allocation of computational resources.
2.2 Downstream Task Performance Scaling
Predicting downstream task performance remains a critical challenge due to emergent abilities in LLMs that manifest only after exceeding task-specific thresholds [37, 28]. Recent works, such as using loss as a proxy [6] or increasing metric resolution [19], have demonstrated potential but encounter challenges in aligning surrogate metrics with original task objectives. Here, we briefly review the two main types of methods for predicting downstream performance:
1. Loss-intermediate prediction. These methods predict the final training loss (or in-domain validation loss) of LLMs with loss scaling laws first, and then predict downstream performance through loss-performance relationships [6, 13, 8]. While these methods leverage established scaling laws for loss predictions, they encounter a fundamental limitation: the inconsistent mapping between loss and performance metrics. In addition, Xiao et al. [38] employ the evaluation set answer loss as an intermediate variable for estimation. Although answer loss correlates with the final performance metrics, its predictability remains low as predicting answer loss shares the challenges with predicting performance, including emergence phenomenon and high variance in task difficulty.
2. End-to-end performance-compute prediction. These methods [19, 26, 1] directly model the relationship between performance and compute (or the number of model parameters). Additionally, Achiam et al. [1] estimate and fit this relationship using a subset of the evaluation set. Hu et al. [19] address the challenge of non-emergent capabilities in smaller models by employing multiple non-greedy decoding evaluations, thereby enabling accurate extrapolation of performance predictions for models with up to 2.4B parameters.
3 Pilot Study
Figure 1: Performance-loss relationship across difference model size (left) and learning rate schedule (middle). At equivalent loss values, smaller models or those with lower learning rates generally achieve higher accuracy than larger models or those with higher learning rates. Performance-compute relationship for different clusters of BBH samples(right). Different clusters demonstrate diverse scaling patterns.
In this section, we present the pilot experiments to illustrate the shortcomings of existing approaches.
Training loss may mismatch downstream tasks performance. Predicting downstream performance based on training loss relies on the assumption that LLMs achieve identical downstream performance at the same loss value–an assumption that often does not hold. In practice, training loss primarily serves as an indicator of in-domain fitting, whereas downstream tasks typically represent out-of-domain evaluations. Moreover, training configurations, such as model size and learning rate, can significantly affect not only the final loss but also the model’s generalization capabilities.
Fig. 1(left) illustrates the performance–loss relationships for LLMs of different sizes on the CEval benchmark [29]. At the same training loss level, smaller models can outperform larger ones in terms of test accuracy. Because smaller models initially exhibit weaker in-domain fitting capacity, they typically require more training steps to reach the same loss value, which can lead to better in-domain generalization once they do. Fig. 1(middle) compares the performance of LLMs trained under different learning rate schedules on the GSM8k dataset [7]. At the same loss level, the performance under the cosine schedule is always worse than that under the constant schedule, indicating that a lower learning rate may prioritize memorization over generalization, thereby diminishing downstream performance.
Diverse scaling patterns within the evaluation set. Scaling patterns capture the performance–compute relationship for a single task sample. However, different task samples exhibit unique computational thresholds, learning slopes, and upper bounds, making it challenging to find a single fitting function (or set of fitting functions) that generalizes well across diverse task samples. Fig. 1(right) illustrates the performance-compute relationships on three random clusters of the BBH benchmark [32], with each cluster containing samples with similar difficulty. Even within a single evaluation set, these scaling curves can vary significantly, indicating that a one-size-fits-all performance-compute curve is insufficient for capturing the full spectrum of a downstream benchmark.
Taken together, these observations highlight the importance of modeling the heterogeneous scaling properties within an evaluation set and identifying a robust intermediate metric to serve as a reliable indicator of the downstream performance of LLMs.
Figure 2: The pipeline of Cluster-On-Difficulty downstream task performance scaling, including 4 stages: a. Represent task difficulty feature with task-wise passrate vector. Cluster on difficulty feature and filter outliers. b. Fit cluster-wise performance-compute curve. Classify clusters into extrapolatable clusters, non-extrapolatable clusters, and non-emergent cluster. c. Predict accuracy on extrapolatable clusters. d. Map subset accuracy prediction to full evaluation set performance.
4 Method
In this section, we introduce the COD method in four parts, illustrated in Fig. 2: 1) We show the advantages of COD and present an improved mean-shift clustering algorithm (Section 4.1); 2) We derive a performance scaling law corresponding to task difficulty variance, which enhances the benefit of extrapolating the performance-compute relationship for task clusters with similar difficulty features (Section 4.2). We fit cluster-wise performance-compute curves on small models and filter extrapolatable clusters; 3) We extrapolate the performance on extrapolatable clusters and predict the accuracy of the target large model on the predictable subset(Section 4.3); 4) We show how to map accuracy on the predictable subset to full evaluations (Section 4.4).
4.1 Clustering on Difficulty
Despite sharing common themes, tasks within evaluation sets demonstrate substantial difficulty differences. These differences result in diverse performance scaling patterns across tasks, making it challenging to apply a universal fitting function for predictions. Instead, we propose clustering tasks with comparable performance scaling behaviors to enable more accurate predictions. This approach minimizes the heterogeneity of difficulty features within clusters while ensuring that each cluster contains a sufficient number of samples for robust evaluation.
We adopt the passrate metric to quantify the capabilities of small-scale models [19]. For each model, we conduct multiple evaluation runs (e.g., 100 trials) and calculate the mean accuracy as the expected probability of correct responses. For each task, we characterize its difficulty through the passrates of models of increasing size. These passrates are arranged in ascending order of model scale, forming feature vectors that ideally exhibit monotonic growth within the [0, 1] range, as model capability typically increases with size. However, we observe that some tasks deviate from the expected scaling pattern, showing non-monotonic difficulty features. This phenomenon may be attributed to metric instability or fluctuations in model performance during training.
Improved clustering methods. We hope to adopt clustering algorithms with the following features: 1. Minimizing intra-class variance to ensure similar extrapolation properties within each cluster, 2. Automatic determination of cluster numbers, as the optimal number varies across evaluation sets and is difficult to pre-specify.
Among classical clustering algorithms, the K-Means algorithm [25] needs to specify the number of clusters in advance. Although there are methods for automatically selecting the optimal number of clusters, e.g. Elbow method [35] and Silhouette [27], these methods need to introduce additional hyperparameters. DBSCAN [11] is a non-parametric density-based clustering algorithm that marks points in low-density regions as outliers while cluster points in the connected high-density regions. In practice, DBSCAN may lead to a larger final intra-class variance and does not meet the clustering requirements of the current task. MeanShift [12] algorithm adopts an extra clustering radius parameter to constrain the intra-class variance, which better fits our demands.
Figure 3: t-SNE visualization of different clustering methods. Each point represents an evaluation sample. DBSCAN(left): Continuous diffusion of clustering leads to too many samples within a class and large inner-group variance. MeanShift(Middle): Keep the high-density region as a unique group. Improved-MeanShift(Right): Constrain inner-group variance with the radius parameter.
To further reduce intro-class variance, we propose an improved MeanShift algorithm to constrain the cluster diameter. At the same time, we maintain a minimum number of tasks in each cluster to reduce metric fluctuations. We provide the t-SNE visualization of evaluation tasks on BBH [32]. Each point represents an evaluation sample and its color denotes the cluster type. Fig. 3. The improved MeanShift prevails, as it effectively splits dense areas into reasonable clusters. We explain the details of clustering algorithms in Section A.1, and smoothing techniques in Section A.2.
4.2 Fitting
Following cluster analysis, we compute evaluation metrics of small models within each cluster and conduct separate extrapolation curve fitting procedures. Small models are trained with the same ratio of training tokens to Compute Per Token (CPT). We propose a scaling law for downstream task performance, supported by theoretical analysis, which allows us to derive prediction formulas for performance scaling within clusters of tasks that share similar difficulty features. The fitting process initially excludes outlier samples, focusing only on the clustered sample set. For each cluster identified in the previous step, we compute accuracy metrics across small models, yielding an expected accuracy array for each cluster. By fitting these accuracy values against the computational costs of small models, we derive the expected accuracy-to-compute curve for each cluster.
We derive the fitting formula for the downstream task scaling law based on the following three assumptions:
  1. 1.
    The relationship between the answer loss and the compute follows a power law, which generalizes the power law in loss prediction into (Question, Answer) format data.
  2. 2.
    For task samples with a finite set of answers, the model gives a random guess choice if it cannot accurately solve it.
  3. 3.
    The task passrate is defined as the product of the predicted probabilities for each token, implying that each task sample has a unique answer, and the model outputs the answer only without any intermediate reasoning progress.
Note that these assumptions may not perfectly hold in practice, we provide additional discussions on Assumption 3 in Section 6. Under the above assumptions we can derive the scaling law for downstream task performance.
Proposition 1 (Scaling Law for Downstream Task Performance).
Given a language model trained with computational budget $C$, and a set of downstream tasks $P$, under the following assumptions: The expected accuracy on tasks $P$ can be modeled as:
$\mathbb{E}_{p}[\mathrm{Acc}(C)]=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{\mu}\right)+o(\mu)$
where:
  • $g$ represents the random guess performance floor;
  • $1-g$ represents the maximum achievable performance improvement;
  • $a,b,c$ are positive constants;
  • $\mu=\frac{1}{\#P}\sum_{(q,\mathrm{ans})\in P}\mathrm{loss}_{\mathrm{ans}}$;
  • $\sigma^{2}=\frac{1}{\#P}\sum(\mathrm{loss}_{\mathrm{ans}\_t}-\mu)^{2}$.
We outline the key proof intuition here, with detailed proofs provided in Appendix B. Like existing approaches [19], we aim to establish the relationship between $\mathrm{loss}_{\mathrm{ans}}$ and model $passrate$, leveraging loss power-law scaling to derive a scaling formula for downstream task passrate metrics.
Proof intuition
The assumption 3 ensures unique task answers and neglecting the impact of model thinking before answers, the probability of correct task completion equals the product of token probabilities in the model output. This implies a negative logarithmic relationship between $\mathrm{loss}_{\mathrm{ans}}$ and $passrate$ for individual tasks.
Previous works overlook that computing the $passrate$ metric for an evaluation set requires averaging $\exp(-\mathrm{loss_{\mathrm{ans}}})$ across tasks, whereas applying the loss scaling law necessitates averaging loss before exponentiation. Mathematically, the performance scaling law computes the arithmetic mean of $\exp(-\mathrm{loss_{\mathrm{ans}}})$, while the loss scaling law after exponentiation yields the geometric mean.
We show that the difference between arithmetic and geometric means can be estimated by $\sigma^{2}/2\mu$, where $\sigma^{2}$ and $\mu$ denote the variance and mean of task passrates, respectively. Consequently, the downstream tasks performance scaling law derived from the loss scaling law is valid only for evaluation sets with limited difficulty variance. Our proposed clustering-based COD method constrains the variance of difficulty features within clusters, enabling better alignment with the performance scaling law.
Finally, we constrain the model output space to a finite answer set, random guessing yields an expected score $g$ for unsuccessful attempts.
Proposition 1 demonstrates that a metric of an evaluation set with similar difficulty features can be effectively modeled using the following formula:
$y(C)=g+(1-g)*e^{-aC^{-b}-c},$ (1)
where $a$ and $b$ jointly influence how accuracy varies with $C$, $c$ controls the upper bound of the fitting curve, and $g$ represents the expected random guess metric for the model on this cluster. Parameters $a$, $b$, $c$, and $g$ are to be fitted.
4.3 Extrapolation
We aim to identify clusters exhibiting robust scaling patterns for reliable performance extrapolation since some clusters have saturated or non-emergent performance on small models and are not expected to give reasonable predictions. We will show that performance prediction on these scalable clusters contributes to the prediction on the full evaluation set. We give the following definition to check whether a cluster is extrapolatable.
Definition: A cluster demonstrates scaling patterns if: (1) its expected accuracy increases monotonically with model size, and (2) the probability of correct responses converges to at least $P$ as computational resources approach infinity, where $P\leq 1$ is a predefined threshold accounting for practical limitations such as ambiguous questions and finite training coverage.
We filter clusters lacking scaling patterns by the following two rules:
  1. 1.
    Negligible accuracy growth with increased computational resources, manifested as minimal $a$ or $b$ values in Eq. 1;
  2. 2.
    Poor extrapolation reliability, indicated by excessive $c$ values in Eq. 1.
In practice, we set the parameter ranges a priori as $a>1$, $b>0.1$, and $10<c<$ 1.
The predictable subset comprises all samples from clusters that exhibit scaling patterns. The final performance prediction for a target model in the predictable subset is computed as the weighted average of the individual cluster predictions, with weights proportional to the cluster sizes.
4.4 Mapping from Predictable Subset to Target Evaluation Set
We extend our predictions from the predictable subset to the complete evaluation set through a principled mapping approach. Our method rests on the observation that extrapolatable and non-extrapolatable samples share question types but differ primarily in difficulty features, suggesting a preserved partial order of metrics across these subsets. We formalize this relationship through a mapping function $f:f(T^{\prime})\rightarrow T$ from predictable subset metrics $T^{\prime}$ to total evaluation set metrics $T$. This function exhibits key properties: 1. continuity and smoothness over $[0,1]$, 2. monotonic increase, and 3. passage through points $(0,0)$ and $(1,1)$. Empirical validation reveals that a quartic function optimally captures this relationship:
$f(x)=\alpha_{1}x^{4}+\alpha_{2}x^{3}+\alpha_{3}x^{2}+(1-\alpha_{1}-\alpha_{2}-\alpha_{3})x$ (2)
To ensure reliable extrapolation, we calibrate the mapping curve using evaluation results of existing models as anchors. Our results show that the subset-to-full mapping generally maintains robustness across model architectures and training data, enabling the use of external models (e.g., Qwen2-72B [39]) as anchors for most tasks. We conduct corresponding experiments in Section 5.5. For data-sensitive tasks, models with similar training distributions provide more reliable anchors, indicating that data consistency takes precedence over architectural variation ensuring mapping accuracy. This calibration strategy enables accurate metric predictions for the complete evaluation set while maintaining computational efficiency.
Finally, combining Eq. 1 and Eq. 2, we get our final metric prediction $p=f\circ y(C_{0})$, where $C_{0}$ is the estimated computation of training the target LLM.
5 Experiments
5.1 Experimental Setups
In our experimental setup, we train several smaller versions of the target model architecture for prediction. These models vary in size but share similar training procedures, with the training data scaled proportionally to their sizes.
Downstream evaluation sets. We adopt the following widely-used benchmarks as our target downstream tasks: For evaluation, we adopt the following widely-used benchmarks, shown in Table 1. Evaluation sets cover popular downstream tasks of the language model, including math, logic, coding, reading comprehension, professional knowledge, etc.
Table 1: Information of evaluation datasets used in the study.
Dataset GSM8K MATH BBH TriviaQA MBPP AGIEval DROP MMLU-pro
[7] [15] [32] [20] [3] [43] [9] [36]
Domain Math Math Reasoning Knowledge Coding Comprehensive Reading Comprehensive
#Questions 1,319 5,000 6,511 17,944 500 8,063 9,536 12,032
#Shots in Prompt 8 4 3 5 3 5 3 5
All models are evaluated in a few-shot in-context learning manner, where they need to generate final answer labels based on given demonstrations and test inputs. We aligned our evaluation setups with LLaMa3 [10].
Model training. To establish performance predictions for large language models, we conduct systematic experiments with a suite of smaller-scale models across different parameter counts and training data volumes, while controlling for other training configurations such as learning rate, batch size, and additional hyperparameters. All models are trained on a constant learning rate scheduler and data with the same distribution. We list model configurations in Table 2.
Table 2: Model architecture specifications across different sizes.
122M 238M 411M 652M 973M 1.9B 7B 12B 70B (Target)
Param. (M) 122 238 411 652 973 1,901 6,980 12,022 68,452
Compute Per Token (B) 1.535 2.684 4.275 6.378 9.060 16.436 54.761 91.609 475.131
Tokens (B) 26 45 72 108 153 277 923 1,544 8,012
Layers 8 10 12 14 16 20 32 43 80
Model Dimension 1,024 1,280 1,536 1,792 2,048 2,560 4,096 4,608 8,192
FFN Dimension 3,584 4,480 5,376 6,272 7,168 8,960 14,336 16,128 28,672
Heads 8 10 12 14 16 20 32 36 64
KV Heads 8 10 12 14 16 20 8 12 8
5.2 Prediction Experiments
Baselines. We evaluate our proposed COD performance scaling for LLMs against existing approaches. The evaluation is conducted on multiple public benchmarks mentioned above, where we utilize a series of smaller models with identical data distribution and architecture but different configurations to estimate the downstream tasks performance of the target 70B large language model.
We compare against three existing prediction methods:
  1. 1.
    End-to-end performance-compute prediction: Extrapolate larger model metrics directly from smaller model evaluation set metrics using performance scaling laws.
  2. 2.
    Passrate-compute prediction: Estimate large model passrates from smaller model passrates [1, 19]. We repeat and evaluate 100 trials for each evaluation set to enhance the performance reliability on smaller models. For a fair comparison, we report the absolution prediction error on the passrate metric instead of greedy decoding accuracy.
  3. 3.
    Loss-intermediate performance prediction: First predict the final training loss of large language model, then estimate downstream task metrics based on the relationship between smaller model evaluation metrics and their corresponding losses [6].
We design two experimental groups to validate the benefits of clustering and the complete pipeline, respectively:
  1. 1.
    COD w/o. mapping: Performing difficulty-based clustering using K-Means, extrapolating within each cluster independently, and then aggregating metrics across clusters without requiring subset-to-full mappings.
  2. 2.
    COD complete: Complete multi-stage proposed approach consisting of clustering, predictable cluster filtering, subset extrapolation, and subset-to-full mapping.
The comparative results across different benchmarks and estimation approaches are presented in Table 3. We evaluate prediction accuracy with the absolute error between predicted and actual performance. We report the prediction error on each single evaluation set and list the mean and the max prediction error.
Table 3: Absolute prediction error on several evaluation sets. A prediction error less than 2 pp is considered an accurate estimate (marked in green), while an error greater than 5 pp is regarded as an invalid estimate (marked in red).
Method Overall Metrics Individual Task Sets
[…]
End-to-end 3.10 6.00 4.00 3.86 0.64 0.68 1.75 6.00 4.11 3.72
Passrate 5.02 8.80 6.71 8.80 3.51 4.00 7.34 6.78 0.26 2.74
Loss-intermediate 5.29 9.39 9.39 6.95 2.33 5.81 5.52 1.41 5.37 5.55
COD (w/o mapping) 2.24 5.26 4.70 0.50 2.91 1.98 0.89 5.26 1.08 0.57
COD (Complete) 1.63 2.38 2.23 1.28 1.77 1.64 2.19 2.38 0.23 1.35
Results. Predictions with an absolute error of less than 2 percentage points (pp) are considered accurate estimations, and when the predicted values fall within the training metric fluctuation range, they are marked in green; predictions with an absolute error greater than 5 indicate invalid estimations, which reduce the overall reliability of the prediction method and are marked in red. These results show our approach significantly outperforms existing methods in both mean and maximum prediction errors, maintaining mean prediction error within 2 pp, thus offering practical guidance for large model training. While existing methods demonstrate good performance on certain evaluation sets, they consistently exhibit substantial estimation errors on a minority of sets, undermining the credibility of their predictions.
Figure 4: Performance-compute relationship across difference prediction method
Through visualization of the performance-compute relationship, we illustrate the distinctive characteristics of different prediction methods, shown in Fig. 4. On the BBH evaluation set, while all three methods yield comparable estimates, end-to-end and loss-intermediate methods demonstrate inadequate fitting for small model evaluation points. In contrast, the cluster method reveals a more sophisticated and well-fitted multi-phase trajectory. For MATH and MMLU pro evaluation sets, the critical challenge involves determining whether large model metrics will experience accelerated growth with increased computing power or encounter performance plateaus. The loss-intermediate method exhibits an underestimation of the model capability ceiling, while the end-to-end method shows prediction errors exceeding 3 pp.
The clustering method’s effectiveness can be attributed to its comprehensive analysis of evaluation set difficulty distributions and scaling laws. It successfully predicts growth patterns in mathematical evaluation sets where most problems demonstrate expanded improvement potential, while accurately capturing the diminishing scaling properties in evaluation sets with score saturation as computational resources increase.
5.3 Comparison of Clustering Methods
We evaluate the impact of clustering methods on the estimation approach. Our goal is to control the average distance between samples and cluster centers within clusters, making difficulty features more similar within clusters. We also ensure that the minimum number of questions in any cluster is not less than 10, considering that too small clusters may lead to instability in metric values. We compared our proposed Improved-MeanShift algorithm with clustering methods including DBScan, MeanShift, and K-Means. Since standard K-Means lacks the ability to filter outliers and directly control intra-cluster distances, we made the following adjustments: (1) Search for the number of clusters such that the minimum cluster size is close to but not less than 10 samples; (2) Draw spheres around cluster centers with a given threshold radius, and treat samples not covered by any sphere as outliers. If a cluster drops to less than 10 samples, we treat its samples as outliers.
We use Intra-cluster Average Distance (IAD) and Outlier Rate (OR) as direct evaluation metrics. With similar OR, a smaller IAD indicates better clustering performance, as shown in Table 4. Additionally, we measure the benefits of different clustering methods on the prediction process by comparing the Extrapolation Errors(EE) of the predictable subset and Final prediction Errors (FE) after clustering, as shown in Table 5.
Table 4: Clustering performance on popular benchmarks.
Method MMLU-pro GSM8k MATH BBH
[…]
K-Means 0.3236 - 0.2238 - 0.2238 - 0.6284 -
DBScan 0.4242 0.56 0.5131 0.53 0.4775 0.68 0.7113 18.92
MeanShift 0.2859 0.39 0.2852 0.61 0.2110 1.44 0.2679 20.72
Improved-KMeans 0.1609 2.85 0.1321 2.73 0.0902 2.22 0.1953 37.23
Improved-MeanShift 0.2225 4.40 0.1854 4.93 0.1463 2.66 0.2143 33.58
Table 5: Prediction errors across clustering algorithms.
Method Mean Max MMLU-pro GSM8k MATH BBH
[…]
K-Means 3.62 3.76 8.16 8.99 3.69 3.69 0.01 0.00 2.62 2.34 8.16 8.99
DBScan 3.93 4.08 4.38 4.36 3.72 3.69 4.08 4.12 4.38 4.16 3.53 4.36
MeanShift 2.12 1.68 3.15 3.08 3.15 3.08 0.67 0.74 2.55 2.26 2.12 0.65
Improved-KMeans 1.33 1.84 3.92 4.08 0.56 0.61 3.92 4.08 0.81 0.51 0.02 2.17
Improved-MeanShift 1.23 1.66 2.20 2.23 1.27 1.35 2.20 2.23 1.14 1.28 0.29 1.77
Table 4 shows that Improved-KMeans and Improved-MeanShift achieve better clustering performance, which is attributed to their incorporation of intra-cluster distance constraints during the clustering process. From Table 5, we can observe that the two methods with better clustering performance correspond to smaller extrapolation errors of estimable subsets and final metric prediction errors.
Although Improved-KMeans achieves optimal clustering performance, its downstream estimation performance on GSM8k is notably inferior compared to other evaluation sets. We believe this is because the K-Means algorithm requires pre-specifying a more explicit number of clusters, but the number of clusters in evaluation sets is difficult to know in advance. While the number of clusters obtained through our search method is effective for some evaluation sets, it lacks stability and ultimately leads to excessive prediction errors in a few evaluation sets. In contrast, our adopted Improved-MeanShift algorithm inherently does not require pre-specifying the number of clusters; instead, it automatically determines this based on our intra-cluster distance constraints. This results in a more stable clustering performance and yields the smallest maximum estimation error across evaluation sets. We present additional clustering experimental results on more evaluation sets in Section C.1, where the conclusions are consistent with existing evaluation sets.
5.4 Extrapolation Formula
To evaluate the effectiveness of different fitting formulas, we conducted an ablation study comparing various formulations of the accuracy-compute relationship. Our baseline formula incorporates random guess probability, exponential decay, and a constant offset term:
$f(C)=g+(1-g)*e^{-aC^{-b}-c}$ (3)
To understand the contribution of each component, we perform ablation experiments by removing or modifying different terms. 1) Without random guess component: $f_{1}(C)=e^{-aC^{-b}-c}$; 2) Without constant term $c$: $f_{2}(C)=g+(1-g)*e^{-aC^{-b}}$; 3) Direct power law relationship [19]: $f_{3}(C)=e^{-aC^{-b}}$.
The comparative results of these formulations are shown in Table 6. Results from three evaluation sets, BBH, Math, and MMLU-pro, are presented here, showing the Extrapolation Error of extrapolatable clusters (EE), the Task Ratio of predictable subset (TR), and the Final prediction Error (FE). These results show that the proposed formula $f$ consistently achieves the smallest extrapolation error and final prediction error, while the ratio of estimable subsets remains similar across different clustering methods.
Table 6: Ablation study results across different benchmarks.
Method BBH MATH MMLU-pro
[…]
Direct Power Law 8.90 49.06 8.88 3.81 81.46 3.35 4.30 95.15 4.27
w/o Random Guess 10.27 45.75 11.20 4.04 81.46 3.55 4.40 95.05 4.37
w/o Constant c 2.14 57.26 4.01 1.40 81.46 1.56 3.85 95.60 3.88
Ours 0.29 52.46 1.77 1.14 81.24 1.28 1.27 94.38 1.35
In the control group, $f_{1}$ performs poorly on tasks with finite answer sets, where small models achieve non-zero scores that $f_{1}$ cannot effectively fit. $f_{2}$ removes $c$, which determines the maximum value of the prediction curve, assuming that the evaluation set performance would reach perfect scores given sufficient computation and parameters. However, this assumption is unreasonable due to the limited training data distribution and ambiguous answers in evaluation set questions, leading to inaccurate predictions. Direct power-law fitting $f_{3}$ fails to model both the metric range constraint of 0 to 1 and the characteristic that metric improvement is typically more difficult near Random Guess capability and capability saturation compared to other regions.
We also observe that TR has little influence on prediction error, which also indicate the robustness of the proposed method. Some evaluation set with low TR due to the non-emergent subset, and our results show that this subset can largely be predicted with metrics of the extrapolatable clusters. We visualize the difficulty distribution of predictable subset and the full evaluation set in Appendix D.
5.5 Anchor Point in Interpolation Mapping
During the mapping phase from estimable subset metrics to full evaluation set metrics, we discovered that models with different training data and architectures exhibit similar mapping relationships. This allows us to leverage metrics from pre-trained models to refine the mapping relationship, thereby improving the accuracy of our final metric estimation.
We use both Qwen2-72B [39] and an in-house model $M$ with a Mixture-of-Experts(MoE) [22] structure trained on the same data distribution as anchor points in the mapping phase. We first obtain the interpolation curve using only small model metrics with points (0,0) and (1,1), then verify the compatibility of anchor points with the existing interpolation curve. When the score of the full set is 0 (1), the score of the subset must also be 0 (1). Results show that despite differences in computational requirements, model architectures, and training data, these models share similar mapping relationships.
This finding also indicates that estimable subset metrics are highly correlated with full-set metrics. Compared to loss-intermediate estimation, estimable subset metrics maintain predictability while reducing interference from other model parameters. Based on these observations, we incorporated the mapping relationships from pre-trained models into the interpolation process, thereby improving both the accuracy and confidence of our large model metric estimations.
We establish three experimental configurations of our method:
  • COD w/o. anchor: The complete estimation process is employed except for not using anchor point interpolation in the Mapping phase.
  • COD w. out-of-distribution (OOD) anchor: The complete proposed methodology incorporates both difficulty-based clustering and predictable subset identification. Using the 72B Qwen2 pretraining model as the anchor model [39].
  • COD w. in-domain(ID) anchor: Using an in-house MoE model with consistent training distribution but different model architecture as the anchor point.
We present the results in Table 7, demonstrating that incorporating both out-of-distribution models and in-distribution models as anchors consistently enhance prediction accuracy. These findings suggest a relatively stable correlation between the metrics of predictable subsets and the full dataset, indicating that the relationship between subset and full-set metrics remains consistent across models trained on different data and with varying architectures. This property enables us to leverage evaluation results from existing models to improve the accuracy of metric predictions for large models with new data and structures. Besides, the division of the evaluation set obtained through clustering is an intrinsic property of the evaluation set itself, independent of the model architecture and training data. Therefore, the predictable subset derived from clustering can also be extended to estimate the metrics of new models. Additionally, we conduct the ablation study on the interpolation method in Section C.2, and results indicate that quartic functions are suitable in our setting.
Table 7: Influence of anchor point usage in the mapping stage.
Method Overall Individual Task Sets
Mean Max GSM8k MATH BBH TriviaQA MBPP AGIEval DROP MMLU-pro
w/o. anchor 3.96 8.80 2.17 5.46 5.08 1.68 8.80 2.38 4.44 1.68
w. ood anchor 1.59 2.56 2.22 0.94 1.84 1.04 2.56 1.86 1.18 1.04
w. id anchor 1.63 2.38 2.23 1.28 1.77 1.64 2.19 2.38 0.23 1.35
6 Conclusion and Discussion
In this work, we introduce a novel downstream performance scaling framework including (1) a difficulty-based clustering approach that effectively models the underlying distribution of each evaluation set; (2) a theoretically grounded scaling law for downstream task performance that provides a fitting formula for performance-compute extrapolation; and (3) a systematic methodology for identifying and leveraging predictable subset that provides a robust intermediate metric for accurate full-set performance predictions.
Our framework, while effective for dense transformers, has not been fully explored for cost-efficient MoE models and does not account for the annealing phase in training, where high-quality data can rapidly enhance performance. The COD method requires sufficient test cases and is not suited for multiple-choice tasks, where performance metrics may diverge from true passrates. Additionally, the framework’s theoretical foundation is insufficient for chain-of-thought reasoning, necessitating future adaptations to address these challenges. We provide detailed discussion in Appendix E.
Looking forward, our approach can be further expanded across model architectures, training methods, and evaluation set types, while extending this framework to address chain-of-thought reasoning patterns offer promising avenues for future research.
Appendix A Improvements of Clustering Algorithm
A.1 Improved MeanShift Algorithm
We iteratively apply the MeanShift algorithm with a predefined cluster radius $R$ and a minimum cluster size $K$. In each iteration, for the clustered samples, we examine whether the distance between each sample and its cluster center exceeds R, and relabel those samples that exceed this threshold as unclustered. For clusters containing fewer than K samples, we mark all samples in these clusters as unclustered. At the end of each iteration, we incorporate both the outliers from MeanShift and our marked unclustered samples into the next round of clustering, continuing this process until no further changes occur in sample labels. We present the pseudocode in Algorithm 1.
Algorithm 1 Iterative MeanShift Clustering Algorithm
1:Initialize all labels in the evaluation set to $-1$
2:repeat
3:     Perform MeanShift clustering with radius $R$ on all samples labeled $-1$
4:     Assign new labels to clustered samples
5:     for each newly labeled sample $i$ do
6:         Calculate distance $dist_{i}$ to its cluster center
7:         if $dist_{i}>R$ then
8:              Reset label to $-1$
9:         end if
10:     end for
11:     for each cluster do
12:         if number of samples in cluster $<K$ then
13:              Reset all samples in this cluster to $-1$
14:         end if
15:     end for
16:     Renumber all non-$\{-1\}$ newly labeled samples to avoid overlap with old labels
17:until no label changes
Filtering Zero-performance Samples
In the evaluation set, there may exist a few extremely difficult problems that require sufficient model parameters to emerge. All small models may fail to solve these problems even after 100 evaluation attempts, resulting in difficulty feature vectors of all zeros. We refer to these as zero-performance samples. Their presence leads to two issues:
  1. 1.
    Zero performance on small models does not necessarily indicate zero accuracy on large models. For these samples, we cannot estimate when emergence will occur or predict large model metrics.
  2. 2.
    During clustering, they may be confused with other low-performing but non-zero samples. Including them in the same cluster would lower the expected accuracy of that cluster, leading to inaccurate fitting and extrapolation later.
Therefore, we pre-filter these zero-performance samples before clustering, treating them as outliers that do not participate in the clustering process. This approach eliminates the need to consider their metrics under large models during subsequent extrapolation and prevents disruption to the clustering of normal difficult samples.
A.2 Smoothing Techniques
Horizontal smoothing: adjacent checkpoint smoothing. Metric fluctuations of individual samples in downstream tasks are not solely due to limited sampling. Another potential factor is noise from uneven data distribution in recent training batches. Therefore, in addition to performing 100 evaluations to mitigate sampling variance, we evaluated 100 times on each of the adjacent checkpoints before and after the selected model. We then averaged these accuracy expectation values across three checkpoints, further reducing sampling variance while offsetting noise from uneven training data distribution. This approach also reduces the number of zero-performance samples, further improving clustering and prediction effectiveness.
Vertical smoothing. Each sample’s features represent the expected correct response rate across models of increasing size, forming a partially ordered sequence. However, the Euclidean distance used for measurement does not consider this sequential information. For example, if a cluster center has a feature sequence of $[0,0,0,0.5]$, sample A with $[0,0,0.2,0.5]$ and sample B with $[0.2,0,0,0.5]$, sample A clearly fits the cluster better than sample B, yet their Euclidean distances are identical.
Note that this smoothing method may not be effective for all downstream tasks. Our current observations suggest that for freeform tasks with limited solution spaces (such as multiple choice, ordering, or judgment questions in freeform format), once models learn to answer within the solution space, their random guess metrics on the evaluation set will be non-zero, more significantly affected by recent training batch data and few-shot cases in prompts. In such cases, vertical smoothing is more likely to bring positive benefits.
In our experiment, we only adopt horizontal smoothing, and leave vertical smoothing as an optional selection.
Appendix B Proof of Proposition
We use Proposition B.1 to derive scaling law for downstream task performance (Proposition B.2).
Proposition B.1 (Arithmetic-geometric mean difference).
For any sequence of positive real numbers $\{x_{i}\}_{i=1}^{n}$, let:
  • $\mu_{a}=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ be the arithmetic mean;
  • $\mu_{g}=\prod_{i=1}^{n}x_{i}^{1/n}$ be the geometric mean;
  • $\sigma^{2}=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu)^{2}$ be the variance.
Then the difference between the arithmetic mean and geometric mean can be estimated as:
$\Delta=\mu_{a}-\mu_{g}=\frac{1}{n}\sum_{i=1}^{n}x_{i}-\left(\prod_{i=1}^{n}x_{i}\right)^{\frac{1}{n}}=\frac{\sigma^{2}}{2\mu_{a}}+o(\mu_{a})$ (4)
Proof.
Taking the logarithm of the geometric mean $\mu_{g}$:
$\log(\mu_{g})=\frac{1}{n}\sum_{i=1}^{n}\log x_{i}$ (5)
Using Taylor expansion of $\log x$ around $\mu$:
$\log x=\log\mu+\frac{x-\mu}{\mu}-\frac{(x-\mu)^{2}}{2\mu^{2}}+o\left((x-\mu)^{2}\right)$ (6)
We can simplify:
$\log(GM)$ $=\frac{1}{n}\sum_{i=1}^{n}\log x_{i}$
$=\log\mu+\frac{1}{n}\sum_{i=1}^{n}\left(\frac{(x_{i}-\mu_{a})}{\mu_{a}}-\frac{(x_{i}-\mu_{a})^{2}}{2\mu_{a}^{2}}\right)+o(\mu_{a})$
$=\log\mu+\frac{1}{\mu}\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}-\mu_{a}\right)}_{\text{equal to 0}}+\frac{1}{2\mu_{a}^{2}}\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu_{a})^{2}\right)}_{\sigma^{2}}+o(\mu_{a})$
$=\log\mu-\frac{\sigma^{2}}{2\mu^{2}}+o(\mu_{a})$
Therefore:
$\mu_{a}-\mu_{g}=\mu_{a}\left(1-\exp{\left(-\frac{\sigma^{2}}{2\mu^{2}}\right)}\right)+o(\mu_{a})$ (7)
When $\frac{\sigma^{2}}{2\mu^{2}}$ is small, this can be approximated as:
$\Delta\approx\frac{\sigma^{2}}{2\mu_{a}}$ (8)
Proposition B.2 (Scaling law for downstream task performance).
Given a language model trained with computational budget $C$, and a set of downstream tasks $P$, under the following assumptions:
  1. 1.
    The relationship between the answer loss and compute follows a power law, $\frac{1}{n}\sum_{(q,\mathrm{ans})}\mathrm{loss}_{\mathrm{ans}}(C)\sim aC^{-b}+c$;
  2. 2.
    For tasks with a finite answer set, the model gives a random guess choice if it cannot truly solve it;
  3. 3.
    Task passrate equals the product of the predicted probability of each token $p_{\mathrm{ans}}=\prod_{t\in\mathrm{ans}}p(t)$, which means that each task has unique answer and model output answer only without thinking progress.
The expected accuracy on tasks $P$ can be modeled as:
$\mathbb{E}_{p}[\mathrm{Acc}(C)]=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{2\mu}\right)+o(\mu)$ (9)
where:
  • $g$ represents the random guess performance floor;
  • $a,b,c$ are positive constants;
  • $\mu=\frac{1}{\#P}\sum_{(q,\mathrm{ans})\in P}loss_{\mathrm{ans}}$;
  • $\sigma^{2}=\frac{1}{\#P}\sum(\mathrm{loss}_{\mathrm{ans}\_t}-\mu)^{2}$.
Proof.
We first use assumption 3 to establish the relationship between model passrate and loss on a task.
$-\log(p_{\mathrm{ans}})=-\log\left(\prod_{t\in\mathrm{ans}}p(t)\right)=-\sum_{t\in\mathrm{ans}}\log(p_{t})=\mathrm{loss}_{\mathrm{ans}}$ (10)
Then take the exponential of both sides, and then take the expectation with respect to different tasks in the evaluation set $p=(q,\mathrm{ans})\in P$. We note that both $p_{\mathrm{ans}}$ and $\mathrm{loss}_{\mathrm{ans}}$ are functions of $C$.
$\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]$ $=\mathbb{E}_{p}[\exp(-\mathrm{loss}_{\mathrm{ans}}(C))]$ (11)
$=\frac{1}{n}\sum_{(q,\mathrm{ans})\in P}\exp(-\mathrm{loss}_{\mathrm{ans}}(C)).$ (12)
We can adopt Proposition B.1 to switch from arithmetic mean to geometric mean of $\mathrm{loss}$, and apply the power law assumption 1.
$\frac{1}{n}\sum_{(q,\mathrm{ans}_{t})\in P}\exp(-\mathrm{loss}_{\mathrm{ans}}(C))$ $=\exp{\underbrace{\left(-\frac{1}{n}\sum_{(q,\mathrm{ans}_{t})\in P}\mathrm{loss}_{\mathrm{ans}}(C)\right)}_{\text{use loss scaling law}}}+\frac{\sigma^{2}}{2\mu}+o(\mu)$ (13)
$=\exp{(-aC^{-b}-c)}+\frac{\sigma^{2}}{2\mu}+o(\mu)$ (14)
where $n=\#P$, and $\mu$, $\sigma^{2}$ follow definitions in the proposition.
Finally, we use assumption 2 to align the answer passrate and the accuracy metric. We can adopt the law of total expectation:
$\mathbb{E}_{p}[\mathrm{Acc}(C)]$ $=\mathbb{P}_{p}(\mathrm{correct})\mathbb{E}[\mathrm{Acc}|\mathrm{correct}]+\mathbb{P}_{p}(\mathrm{incorrect})\mathbb{E}[\mathrm{Acc|incorrect}]$ (15)
Note that $\mathbb{P}_{p}(\mathrm{correct})=\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]$, $\mathbb{E}[\mathrm{Acc}|\mathrm{correct}]=1$ and $\mathbb{P}_{p}(\mathrm{incorrect})=1-\mathbb{P}_{p}(\mathrm{correct})$. We also define $g$ as the random guess accuracy performance, thus we have $\mathbb{E}[\mathrm{Acc|incorrect}]=g$. Take these results into Eq. 15, and we have:
$\mathbb{E}_{p}[\mathrm{Acc}(C)]$ $=\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]+(1-\mathbb{E}_{p}[p_{\mathrm{ans}}(C)])*g$ (16)
$=g+(1-g)\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]$ (17)
$=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{2\mu}\right)+o(\mu)$ (18)
Proposition B.2 demonstrates that a metric of an evaluation set with similar difficulty features can be effectively modeled using the following formula:
$f(C)=g+(1-g)\exp{(-aC^{-b}-c)}$ (19)
Appendix C Additional Ablation Studies
C.1 Comparison of Clustering Methods on Extra Evaluation Sets.
We provide additional clustering evaluation results across more evaluation sets in Table C1 and Table C2, which maintain consistency with the conclusions presented in the main text.
Table C1: Clustering performance on advanced task benchmarks (IAD: Intra-cluster Average Distance, OR: Outlier Rate)
Method TriviaQA AGIEval DROP MBPP
[…]
K-Means 0.4388 - 0.4572 - 0.5554 - 0.3383 -
DBScan 0.7039 6.38 0.5591 3.67 0.6651 11.08 0.5060 12.80
MeanShift 0.2521 6.77 0.2886 2.99 0.2507 11.81 0.2167 15.60
Improved-KMeans 0.1239 11.97 0.1536 7.60 0.1428 21.42 0.1667 19.40
Improved-MeanShift 0.1871 11.54 0.2100 11.50 0.1974 19.88 0.1745 21.60
Table C2: Prediction errors on advanced task benchmarks.
Method Mean Max TriviaQA AGIEval DROP MBPP
EE FE EE FE EE FE EE FE EE FE EE FE
K-Means 2.44 3.36 2.97 8.99 2.97 2.46 2.61 2.68 1.66 1.64 2.53 2.67
DBScan 3.04 2.79 6.43 4.36 1.11 0.81 6.43 6.27 3.03 2.66 1.57 1.41
MeanShift 3.21 2.34 4.18 4.90 3.64 4.90 2.63 3.23 4.18 4.00 2.40 1.22
Improved-KMeans 3.38 3.80 5.96 5.56 1.18 1.12 5.96 5.56 3.99 5.24 2.39 3.25
Improved-MeanShift 1.13 1.61 1.58 2.38 1.58 1.64 1.11 2.38 0.26 0.23 1.56 2.19
C.2 Interpolation Method
To evaluate different interpolation methods for prediction accuracy, we compared various mathematical approaches. Our baseline method uses quartic polynomial interpolation, which we compare against several alternative approaches, including Cubic spline interpolation, Cubic polynomial interpolation, and Quintic polynomial interpolation
The comparative results across different benchmarks are shown in Table C3. We report the prediction error between the real performance of a large model and the mapping result.
Table C3: Comparison of different interpolation methods across benchmarks.
Prediction Error BBH Math MMLU-pro
Cubic Spline 0.68 1.31 1.37
Cubic Polynomial 3.38 1.12 1.35
Quintic Polynomial 0.18 1.42 1.36
Quartic Polynomial 1.77 1.28 1.35
Furthermore, in Fig. C1, we plot the mapping process using different interpolation formulas, where the x-axis represents the predictable subset indices and the y-axis represents the full set indices. The red points are the numerical values to be fitted, green points represent predicted values, purple points represent anchor points and blue points show the actual performance.
The prediction performance shows certain robustness across different interpolation methods. We aim to use the simplest possible interpolation function while maintaining low prediction errors. Based on the above results, we observe that the Cubic Polynomial shows larger prediction errors due to underfitting on BBH, while the Cubic Spline exhibits some overfitting. Both Quartic Polynomial and Quintic Polynomial perform well, therefore we chose the Quartic Polynomial method as it requires fewer fitting parameters.
Figure C1: Performance mapping with different interpolation methods on BBH evaluation set. The cubic spline is overfitted, and the cubic polynomial method is underfitted. Quartic polynomials and quintic polynomials are comparable while quartic polynomial has fewer parameters.
Appendix D Difficulty Distribution of Predictable Subset
We analyze the proportion of predictable subset tasks across different difficulty levels. The difficulty distributions of predictable subset versus complete sets for different evaluation benchmarks are illustrated in Fig. D2. We use the scores from the 12B model as the basis for difficulty classification. The results show that MMLU-pro and GSM8k evaluation sets have larger proportions of predictable subset, indicating that most questions in these datasets exhibit good performance scaling properties. In contrast, many difficult questions with near-zero scores in the Math evaluation set fall outside the predictable subset, requiring adjustment during the mapping phase. Meanwhile, BBH shows consistent proportions of predictable subset across difficulty levels, as some of its questions demonstrate oscillatory patterns with limited improvement despite increased computing.
The proportion of predictable subset can serve as a metric for assessing evaluation set quality. Evaluation sets with larger predictable subset yield more reliable experimental conclusions from smaller models. When constructing evaluation sets, we recommend screening or supplementing unpredictable clusters and ensuring a minimum number of questions for each difficulty feature to reduce metric volatility.
Figure D2: Difficulty distribution comparison on a 12B model between predictable subset and full evaluation set.
Appendix E Limitations
Influence of model structure and training configurations. Mixture-of-Experts (MoE) models excel in training and inference cost, and are widely used in production. In this work, we reveal the performance scaling on dense transformers, while prediction on MoE models is still underexplored. However, we believe that the proposed method is not significantly affected by the model architecture. If we apply the complete pipeline to MoE models, we expect to achieve similar results. Besides, we only study the pre-training performance prediction with a constant learning rate and do not cover the impact of the annealing training. In this phase, higher-quality data is usually adopted, which can rapidly improve the model’s capabilities. As a result, performance prediction faces greater challenges.
Category of evaluation sets. The proposed Clustering-on-Difficulty method requires a sufficient number of test cases, as too few samples can lead to unstable cluster metrics and ineffective estimation. However, from an evaluation set design perspective, an evaluation set with good predictive properties enables more effective generalization from small-scale to large-scale models, thus providing better guidance for model iteration.
Furthermore, for multiple-choice tasks, the model only needs to assign a higher probability to the correct option compared to others, creating a discrepancy between this metric and the model’s true passrate. Given that more evaluation sets are adopting the Chain-of-Thoughts prompts, we have not included multiple-choice tasks that only require option selection.
Chain-of-thought performance prediction. Proposition B.2 assumes that evaluation sets directly assess models’ ability to provide answers. However, increasingly more evaluations allow models to think before providing answers. Recent works on inference time scaling [31, 4] further demonstrate that for tasks involving mathematics, reasoning, and coding, training models to complete tasks through longer inference computation can significantly improve downstream task performance. In cases where the reasoning process or answers are not unique, the relationship between a model’s answer loss and passrate on a task may not necessarily follow the exponential relationship between the answer loss and the sample passrate. Although our approach maintains its prediction effectiveness in such situations, the theoretical explanation for these cases is insufficient. Therefore, we consider improving prediction methods based on chain-of-thought characteristics and expanding theoretical foundations as future work.

### v2

@@
 Abstract
-The rapid advancements in computing dramatically increase the scale and cost of training Large Language Models (LLMs).
-Accurately predicting downstream task performance prior to model training is crucial for efficient resource allocation, yet remains challenging due to two primary constraints: (1) the “emergence phenomenon”, wherein downstream performance metrics become meaningful only after extensive training, which limits the ability to use smaller models for prediction; (2) Uneven task difficulty distributions and the absence of consistent scaling laws, resulting in substantial metric variability.
-Existing performance prediction methods suffer from limited accuracy and reliability, thereby impeding the assessment of potential LLM capabilities.
-To address these challenges, we propose a Clustering-On-Difficulty (COD) downstream performance prediction framework.
-COD first constructs a predictable support subset by clustering tasks based on difficulty features, strategically excluding non-emergent and non-scalable clusters.
-The scores on the selected subset serve as effective intermediate predictors of downstream performance on the full evaluation set.
-With theoretical support, we derive a mapping function that transforms performance metrics from the predictable subset to the full evaluation set, thereby ensuring accurate extrapolation of LLM downstream performance.
-The proposed method has been applied to predict performance scaling for a 70B LLM, providing actionable insights for training resource allocation and assisting in monitoring the training process.
-Notably, COD achieves remarkable predictive accuracy on the 70B LLM by leveraging an ensemble of small models, demonstrating an absolute mean deviation of 1.36% across eight important LLM evaluation benchmarks.
+The escalating scale and cost of Large Language Models (LLMs) training necessitate accurate pre-training prediction of downstream task performance for efficient resource allocation.
+This is challenged by: 1) the emergence phenomenon, where metrics become meaningful only after extensive training, hindering prediction by smaller models; and 2) uneven task difficulty and inconsistent performance scaling patterns, leading to high metric variability.
+Current prediction methods lack accuracy and reliability.
+We propose a Clustering-On-Difficulty (COD) framework for downstream performance prediction.
+The COD framework clusters tasks by their difficulty scaling features, thereby establishing a more stable and predictable support subset through the exclusion of tasks exhibiting non-emergent behavior or irregular scaling.
+We adopt a performance scaling law to predict cluster-wise performance with theoretical support.
+Predictable subset performance acts as an intermediate predictor for the full evaluation set.
+We further derive a mapping function to accurately extrapolate the performance of the subset to the full set.
+Applied to an LLM with 70B parameters, COD achieved a 1.36% average prediction error across eight key LLM benchmarks, offering actionable insights for resource allocation and training monitoring of LLMs pretraining.
+[…]
+Chenyin Xu at , Ke Shen at
 1 Introduction
@@
 Specifically, performance scaling of downstream tasks aims to predict the accuracy of the target LLM on downstream tasks using metrics from smaller models.
-Our objective is to develop a prediction method that works reliably across a diverse range of downstream tasks, minimizing the worst-case prediction error.
+Our objective is to develop a prediction method that works reliably on a diverse range of downstream tasks, optimizing the worst-case prediction error.
 Despite extensive efforts, a reliable scaling law for downstream tasks remains elusive.
-One line of work attempts to extrapolate large-model performance by modeling the performance-loss relationship [6, 13, 8, 38, 26], but this often fails to capture the emergent behaviors of LLMs and the mismatch between in-domain loss and downstream metrics [42].
-Another line of research focuses on direct extrapolation of performance-compute relationship [1, 19], yet the uneven difficulty distribution across different evaluation samples undermines its accuracy.
+One line of work attempts to extrapolate the performance of a large model by modeling the performance-loss relationship [6, 13, 8, 38, 28], but this often fails to capture the emergent behaviors of LLMs and the mismatch between the in-domain loss and downstream metrics [43].
+Another line of research focuses on direct extrapolation of the performance-compute relationship [1, 20], yet the uneven difficulty distribution between different evaluation samples undermines its accuracy.
+A critical oversight in current approaches is the assumption of uniform scaling behavior across diverse evaluation samples.
 We observe that different evaluation samples actually follow distinct performance scaling patterns, and thus applying a single extrapolation formula to the entire evaluation set is suboptimal.
-We give the detailed analysis in Section 3.
+We give a detailed analysis in Sec.
+3.
 To address these challenges, we propose a new performance scaling law, derived from the existing loss scaling law [22], specifically applicable to evaluation subsets that exhibit consistent performance scaling patterns.
 Building on the performance scaling law, we develop a Clustering-On-Difficulty (COD) multi-stage framework for predicting downstream performance.
-Specifically, we first cluster tasks by their difficulty features, and then filter out clusters that lack valid extrapolation patterns.
+Specifically, we first filter out clusters that lack scaling properties using an improved MeanShift clustering algorithm.
 Next, we fit the performance-compute relationships in the remaining clusters under our performance scaling law, extrapolate the performance of large models within these clusters, and finally map the aggregated predictions to the complete task set.
-We validate our COD approach on eight evaluation sets, including popular MATH [15], BBH [32], and MMLU pro [36].
-COD achieves an average prediction error of 1.36% on a 70B-parameter LLM.
+We validate our COD approach on eight popular evaluation sets, including MATH [16], BBH [33], and MMLU pro [36] datasets.
+COD achieves an average prediction error of 1.36% on an LLM with 70B parameters.
 Our results demonstrate that this difficulty-aware framework substantially outperforms existing methods, establishing a promising paradigm for accurate downstream performance scaling of LLMs.
-Our contributions can be summarized as follows:
-  • We propose the COD framework to address high variance and emergent phenomena in the LLM performance by effectively modeling the difficulty distribution within the evaluation sets.
+Our contributions can be summarized as follows.
+  • We propose the COD framework to address high variance and emergent phenomena in LLM performance scaling by effectively modeling the difficulty distribution within the evaluation sets.
   • We introduce a performance scaling law for cluster-wise performance prediction, with theoretical support and experimental validation.
-  • Extensive experiments conducted across eight diverse evaluation sets demonstrate that COD achieves a state-of-the-art average prediction error of 1.36% on a 70B-parameter LLM.
+  • Extensive experiments conducted in eight different evaluation sets demonstrate that COD provides reliable predictions with an average prediction error of 1. 36% on an LLM with 70B parameters.
 2 Related Work
 2.1 Loss Scaling Laws
-Loss scaling laws provide a systematic framework for understanding the relationship between computational resources, data, model size, and the final performance of LLMs.
+Loss scaling laws provide a systematic framework for understanding the relationship between computational resources, data, model size, and the LLM performance.
 Early work by Kaplan et al.
 [22] demonstrates that the pre-training loss of LLMs follows a power-law relationship with the compute (the number of floating-point operations) used in training.
-Subsequent studies extend these findings to other domains, such as computer vision [41], graph learning [24] and vision-language models [2, 16].
+Subsequent studies extend these findings to other domains, such as computer vision [42], graph learning [26], and vision-language models [2, 17].
 Recent research has also explored scaling laws in specific contexts, such as fine-tuning [18, 35], vocabulary size optimization [34], retrieval-augmented models [31], and hyperparameter tuning [24, 41].
@@
 Here, we briefly review the two main types of methods for predicting downstream performance:
-1.
 Loss-intermediate prediction.
@@
 Although answer loss correlates with the final performance metrics, its predictability remains low as predicting answer loss shares the challenges with predicting performance, including emergence phenomenon and high variance in task difficulty.
-2.
 End-to-end performance-compute prediction.
@@
 Additionally, Achiam et al.
-[1] estimate and fit this relationship using a subset of the evaluation set.
+[1] estimate and fit this relationship using a subset of the evaluation set, while still failing to predict the full set.
 Hu et al.
@@
 3 Pilot Study
-Figure 1: Performance-loss relationship across difference model size (left) and learning rate schedule (middle).
-At equivalent loss values, smaller models or those with lower learning rates generally achieve higher accuracy than larger models or those with higher learning rates.
-Performance-compute relationship for different clusters of BBH samples(right).
-Different clusters demonstrate diverse scaling patterns.
+Figure 1: Performance-loss relationship across different model sizes (left) and learning rate schedules (middle). Performance-compute relationship for different clusters of the BBH samples(right)
 In this section, we present the pilot experiments to illustrate the shortcomings of existing approaches.
-Training loss may mismatch downstream tasks performance.
-Predicting downstream performance based on training loss relies on the assumption that LLMs achieve identical downstream performance at the same loss value–an assumption that often does not hold.
+Training loss may mismatch downstream task performance.
+Predicting downstream performance from training loss assumes LLMs achieve identical downstream results at the same loss value, which does not hold universally.
 In practice, training loss primarily serves as an indicator of in-domain fitting, whereas downstream tasks typically represent out-of-domain evaluations.
@@
 Diverse scaling patterns within the evaluation set.
-Scaling patterns capture the performance–compute relationship for a single task sample.
-However, different task samples exhibit unique computational thresholds, learning slopes, and upper bounds, making it challenging to find a single fitting function (or set of fitting functions) that generalizes well across diverse task samples.
+Different task samples exhibit unique computational thresholds, learning slopes, and upper bounds, making it challenging to find a single fitting function (or function group) that generalizes well across diverse task samples.
 Fig.
 1(right) illustrates the performance-compute relationships on three random clusters of the BBH benchmark [33], with each cluster containing samples with similar difficulty.
-Even within a single evaluation set, these scaling curves can vary significantly, indicating that a one-size-fits-all performance-compute curve is insufficient for capturing the full spectrum of a downstream benchmark.
+Even within a single evaluation set, these scaling curves can vary significantly, indicating that a one-size-fits-all performance-compute curve is insufficient for capturing the full spectrum of a downstream evaluation set.
 Taken together, these observations highlight the importance of modeling the heterogeneous scaling properties within an evaluation set and identifying a robust intermediate metric to serve as a reliable indicator of the downstream performance of LLMs.
@@
 Represent task difficulty feature with task-wise passrate vector.
-Cluster on difficulty feature and filter outliers.
+Cluster on the difficulty feature and filter outliers.
 b.
 Fit cluster-wise performance-compute curve.
-Classify clusters into extrapolatable clusters, non-extrapolatable clusters, and non-emergent cluster.
+Classify clusters into extrapolatable clusters, non-extrapolatable clusters, and non-emergent clusters.
 c.
@@
 4 Method
-In this section, we introduce the COD method in four parts, illustrated in Fig.
-2: 1) We show the advantages of COD and present an improved mean-shift clustering algorithm (Section 4.1); 2) We derive a performance scaling law corresponding to task difficulty variance, which enhances the benefit of extrapolating the performance-compute relationship for task clusters with similar difficulty features (Section 4.2).
-We fit cluster-wise performance-compute curves on small models and filter extrapolatable clusters; 3) We extrapolate the performance on extrapolatable clusters and predict the accuracy of the target large model on the predictable subset(Section 4.3); 4) We show how to map accuracy on the predictable subset to full evaluations (Section 4.4).
+In this section, we introduce the COD method in four parts, as illustrated in Fig.
+2: 1) We show the advantages of COD and present an improved mean-shift clustering algorithm (Sec.
+4.1); 2) We derive a performance scaling law corresponding to task difficulty variance, which supports extrapolating the performance-compute relationship for task clusters with similar difficulty features (Sec.
+4.2).
+We fit cluster-wise performance-compute curves on small models and filter extrapolatable clusters; 3) We extrapolate the performance on extrapolatable clusters and predict the accuracy of the target large model on the predictable subset(Sec.
+4.3); 4) We show how to map accuracy on the predictable subset to full evaluations (Sec.
+4.4).
 4.1 Clustering on Difficulty
-Despite sharing common themes, tasks within evaluation sets demonstrate substantial difficulty differences.
-These differences result in diverse performance scaling patterns across tasks, making it challenging to apply a universal fitting function for predictions.
-Instead, we propose clustering tasks with comparable performance scaling behaviors to enable more accurate predictions.
-This approach minimizes the heterogeneity of difficulty features within clusters while ensuring that each cluster contains a sufficient number of samples for robust evaluation.
-We adopt the passrate metric to quantify the capabilities of small-scale models [19].
-For each model, we conduct multiple evaluation runs (e.g., 100 trials) and calculate the mean accuracy as the expected probability of correct responses.
-For each task, we characterize its difficulty through the passrates of models of increasing size.
-These passrates are arranged in ascending order of model scale, forming feature vectors that ideally exhibit monotonic growth within the [0, 1] range, as model capability typically increases with size.
-However, we observe that some tasks deviate from the expected scaling pattern, showing non-monotonic difficulty features.
-This phenomenon may be attributed to metric instability or fluctuations in model performance during training.
+Tasks within evaluation sets, despite common themes, show substantial difficulty differences, leading to diverse scaling patterns that hinder universal fitting functions.
+We propose clustering tasks by similar performance scaling behaviors to minimize intra-cluster heterogeneity while ensuring robust evaluation with sufficient samples per cluster.
+To quantify task difficulty, we leverage the passrate(mean accuracy across multiple trials) from small models of progressively increasing scales.
+This process generates a feature vector for each task, which ideally exhibits monotonic growth, thereby reflecting the incremental enhancement of model capabilities.
 Improved clustering methods.
@@
 Automatic determination of cluster numbers, as the optimal number varies across evaluation sets and is difficult to pre-specify.
-Among classical clustering algorithms, the K-Means algorithm [25] needs to specify the number of clusters in advance.
-Although there are methods for automatically selecting the optimal number of clusters, e.g.
-Elbow method [35] and Silhouette [27], these methods need to introduce additional hyperparameters.
-DBSCAN [11] is a non-parametric density-based clustering algorithm that marks points in low-density regions as outliers while cluster points in the connected high-density regions.
-In practice, DBSCAN may lead to a larger final intra-class variance and does not meet the clustering requirements of the current task.
-MeanShift [12] algorithm adopts an extra clustering radius parameter to constrain the intra-class variance, which better fits our demands.
-Figure 3: t-SNE visualization of different clustering methods.
-Each point represents an evaluation sample.
-DBSCAN(left): Continuous diffusion of clustering leads to too many samples within a class and large inner-group variance.
-MeanShift(Middle): Keep the high-density region as a unique group.
-Improved-MeanShift(Right): Constrain inner-group variance with the radius parameter.
+While K-Means [27] requires pre-specifying cluster numbers and DBSCAN [11] may yield large intra-cluster variance, MeanShift [12] allows constraining this variance via a radius parameter, better suiting our needs.
+Figure 3: t-SNE visualization of different clustering methods: DBSCAN(left), MeanShift(Middle), Improved-MeanShift(Right). Each point represents an evaluation sample.
 To further reduce intro-class variance, we propose an improved MeanShift algorithm to constrain the cluster diameter.
@@
 We provide the t-SNE visualization of evaluation tasks on BBH [33].
-Each point represents an evaluation sample and its color denotes the cluster type.
-Fig.
-3.
-The improved MeanShift prevails, as it effectively splits dense areas into reasonable clusters.
-We explain the details of clustering algorithms in Section A.1, and smoothing techniques in Section A.2.
+Each point represents an evaluation sample, and its color denotes the cluster type.
+As shown in Fig.
+3, our improved MeanShift effectively splits dense areas.
+We explain the details of clustering algorithms in Sec.
+A.1, and smoothing techniques in Sec.
+A.2.
 4.2 Fitting
-Following cluster analysis, we compute evaluation metrics of small models within each cluster and conduct separate extrapolation curve fitting procedures.
-Small models are trained with the same ratio of training tokens to Compute Per Token (CPT).
-We propose a scaling law for downstream task performance, supported by theoretical analysis, which allows us to derive prediction formulas for performance scaling within clusters of tasks that share similar difficulty features.
-The fitting process initially excludes outlier samples, focusing only on the clustered sample set.
-For each cluster identified in the previous step, we compute accuracy metrics across small models, yielding an expected accuracy array for each cluster.
-By fitting these accuracy values against the computational costs of small models, we derive the expected accuracy-to-compute curve for each cluster.
-We derive the fitting formula for the downstream task scaling law based on the following three assumptions:
+After clustering, we compute metrics for small models within each cluster.
+These small models are trained with the same ratio of training tokens to compute per token.
+We then apply a novel performance scaling law, based on theoretical analysis, to fit accuracy-to-compute curves for each cluster, focusing on clustered samples after excluding outliers.
+We derive the following fitting formula for the downstream task scaling law based on the training loss scaling law.
+Theorem 1 (Scaling Law for Downstream Task Performance).
+Given a language model trained with computational budget $C$, and a set of downstream tasks $P$, under the following assumptions:
   1. 1.
-    The relationship between the answer loss and the compute follows a power law, which generalizes the power law in loss prediction into (Question, Answer) format data.
+    Given a task $(q,\mathrm{ans})\in P$, the relationship between the answer loss and the compute follows a power law, $\frac{1}{n}\sum_{(q,\mathrm{ans})}\mathrm{loss}_{\mathrm{ans}}(C)\sim aC^{-b}+c$;
   2. 2.
-    For task samples with a finite set of answers, the model gives a random guess choice if it cannot accurately solve it.
+    For tasks with a finite answer set, the model gives a random guess choice if it cannot truly solve it, i.e. $\mathbb{E}[\mathrm{Acc|incorrect}]=g$;
   3. 3.
-    The task passrate is defined as the product of the predicted probabilities for each token, implying that each task sample has a unique answer, and the model outputs the answer only without any intermediate reasoning progress.
-Note that these assumptions may not perfectly hold in practice, we provide additional discussions on Assumption 3 in Section 6.
-Under the above assumptions we can derive the scaling law for downstream task performance.
-Proposition 1 (Scaling Law for Downstream Task Performance).
-Given a language model trained with computational budget $C$, and a set of downstream tasks $P$, under the following assumptions: The expected accuracy on tasks $P$ can be modeled as:
-$\mathbb{E}_{p}[\mathrm{Acc}(C)]=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{\mu}\right)+o(\mu)$
-where:
-  • $g$ represents the random guess performance floor;
-  • $1-g$ represents the maximum achievable performance improvement;
-  • $a,b,c$ are positive constants;
-  • $\mu=\frac{1}{\#P}\sum_{(q,\mathrm{ans})\in P}\mathrm{loss}_{\mathrm{ans}}$;
-  • $\sigma^{2}=\frac{1}{\#P}\sum(\mathrm{loss}_{\mathrm{ans}\_t}-\mu)^{2}$.
-We outline the key proof intuition here, with detailed proofs provided in Appendix B.
-Like existing approaches [19], we aim to establish the relationship between $\mathrm{loss}_{\mathrm{ans}}$ and model $passrate$, leveraging loss power-law scaling to derive a scaling formula for downstream task passrate metrics.
-Proof intuition
-The assumption 3 ensures unique task answers and neglecting the impact of model thinking before answers, the probability of correct task completion equals the product of token probabilities in the model output.
-This implies a negative logarithmic relationship between $\mathrm{loss}_{\mathrm{ans}}$ and $passrate$ for individual tasks.
-Previous works overlook that computing the $passrate$ metric for an evaluation set requires averaging $\exp(-\mathrm{loss_{\mathrm{ans}}})$ across tasks, whereas applying the loss scaling law necessitates averaging loss before exponentiation.
-Mathematically, the performance scaling law computes the arithmetic mean of $\exp(-\mathrm{loss_{\mathrm{ans}}})$, while the loss scaling law after exponentiation yields the geometric mean.
-We show that the difference between arithmetic and geometric means can be estimated by $\sigma^{2}/2\mu$, where $\sigma^{2}$ and $\mu$ denote the variance and mean of task passrates, respectively.
-Consequently, the downstream tasks performance scaling law derived from the loss scaling law is valid only for evaluation sets with limited difficulty variance.
-Our proposed clustering-based COD method constrains the variance of difficulty features within clusters, enabling better alignment with the performance scaling law.
-Finally, we constrain the model output space to a finite answer set, random guessing yields an expected score $g$ for unsuccessful attempts.
-Proposition 1 demonstrates that a metric of an evaluation set with similar difficulty features can be effectively modeled using the following formula:
+    Task passrate equals the product of the predicted probability of each token $p_{\mathrm{ans}}=\prod_{t\in\mathrm{ans}}p(t)$, which means that each task has a unique answer and model output answer only without thinking progress.
+The expected accuracy on tasks $P$ can be modeled as:
+$\mathbb{E}_{p}[\mathrm{Acc}(C)]=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{\mu}\right)+o(\mu),$ (1)
+where $g$ represents the random guess performance floor, $1-g$ represents the maximum achievable performance improvement, $a,b,c$ are positive constants, and $\mu=\frac{1}{\#P}\sum_{(q,\mathrm{ans})\in P}\mathrm{loss}_{\mathrm{ans}}$; $\sigma^{2}=\frac{1}{\#P}\sum(\mathrm{loss}_{\mathrm{ans}_{t}}-\mu)^{2}$.
+Proof Sketch.
+Assumption 3 (unique answers, no pre-answer reasoning) implies a negative logarithmic relationship between $\mathrm{loss}_{\mathrm{ans}}$ and individual task passrate, as correct task completion probability becomes the product of answer token probabilities.
+A key distinction, often overlooked, is how averaging occurs: the passrate metric averages $\exp(-\mathrm{loss}_{\mathrm{ans}})$ (an arithmetic mean), whereas applying the loss scaling law effectively averages loss before exponentiation (yielding a geometric mean).
+We show that the difference between these means can be approximated by $\frac{\sigma^{2}}{\mu}$ (where $\sigma^{2}$ and $\mu$ are the variance and mean of task passrates, respectively).
+Consequently, a performance scaling law derived from the loss scaling law is most valid for evaluation sets with limited difficulty variance.
+Our COD method achieves this by clustering tasks, thereby constraining intra-cluster difficulty variance and enabling better alignment with the performance scaling law.
+Finally, for tasks with a finite answer set, a random guessing score $g$ accounts for expected performance on unmastered tasks.
+Detailed proofs are provided in Appendix B.
+∎
+theorem 1 demonstrates that a metric of an evaluation set with similar difficulty features can be effectively modeled using the following formula:
 $y(C)=g+(1-g)*e^{-aC^{-b}-c},$ (2)
-where $a$ and $b$ jointly influence how accuracy varies with $C$, $c$ controls the upper bound of the fitting curve, and $g$ represents the expected random guess metric for the model on this cluster.
+where $a$ and $b$ jointly influence how accuracy varies with $C$, $c$ controls the upper bound of the fitting curve, and $g$ represents the expected random guess metric for a task cluster.
 Parameters $a$, $b$, $c$, and $g$ are to be fitted.
+Note that these assumptions may not perfectly hold in practice, we provide additional discussions on assumption 3 in Appendix G.
 4.3 Extrapolation
-We aim to identify clusters exhibiting robust scaling patterns for reliable performance extrapolation since some clusters have saturated or non-emergent performance on small models and are not expected to give reasonable predictions.
-We will show that performance prediction on these scalable clusters contributes to the prediction on the full evaluation set.
-We give the following definition to check whether a cluster is extrapolatable.
-Definition: A cluster demonstrates scaling patterns if: (1) its expected accuracy increases monotonically with model size, and (2) the probability of correct responses converges to at least $P$ as computational resources approach infinity, where $P\leq 1$ is a predefined threshold accounting for practical limitations such as ambiguous questions and finite training coverage.
-We filter clusters lacking scaling patterns by the following two rules:
+To ensure reliable extrapolation, we identify clusters exhibiting robust scaling patterns, as some clusters may show saturated or non-emergent performance on smaller models, making them unsuitable for prediction.
+A cluster is deemed extrapolatable if it meets two criteria: (1) its expected accuracy increases monotonically with model size, and (2) its performance converges to at least a predefined threshold $P$ (where $P\leq 1$ accounts for practical limits like ambiguous questions or finite training coverage).
+We filter out non-extrapolatable clusters using two rules based on the parameters from Eq. 2:
   1. 1.
-    Negligible accuracy growth with increased computational resources, manifested as minimal $a$ or $b$ values in Eq. 1;
+    Negligible accuracy growth, indicated by minimal $a$ or $b$ values.
   2. 2.
-    Poor extrapolation reliability, indicated by excessive $c$ values in Eq. 1.
-In practice, we set the parameter ranges a priori as $a>1$, $b>0.1$, and $10<c<$ 1.
-The predictable subset comprises all samples from clusters that exhibit scaling patterns.
-The final performance prediction for a target model in the predictable subset is computed as the weighted average of the individual cluster predictions, with weights proportional to the cluster sizes.
+    Poor extrapolation reliability, indicated by an excessive $c$ value.
+In practice, for extrapolatable clusters, we set $a>1$, $b>0.1$, and $10\leq c<$ 1.
+The clusters satisfying these conditions form the predictable subset.
+The final performance prediction for a target model on this subset is the weighted average of the extrapolated predictions from these individual clusters, with weights proportional to cluster sizes.
 4.4 Mapping from Predictable Subset to Target Evaluation Set
-We extend our predictions from the predictable subset to the complete evaluation set through a principled mapping approach.
-Our method rests on the observation that extrapolatable and non-extrapolatable samples share question types but differ primarily in difficulty features, suggesting a preserved partial order of metrics across these subsets.
-We formalize this relationship through a mapping function $f:f(T^{\prime})\rightarrow T$ from predictable subset metrics $T^{\prime}$ to total evaluation set metrics $T$.
-This function exhibits key properties: 1.
-continuity and smoothness over $[0,1]$, 2.
-monotonic increase, and 3.
-passage through points $(0,0)$ and $(1,1)$.
-Empirical validation reveals that a quartic function optimally captures this relationship:
+We map predictions from the predictable subset $T^{\prime}$ to the complete evaluation set $T$ using a function $f$.
+This approach is based on the observation that extrapolatable and non-extrapolatable samples, while differing in difficulty, often share question types, suggesting a preserved relative metric ordering.
+The mapping function $f:T^{\prime}\rightarrow T$ is continuous, smooth over $[0,1]$, monotonically increasing, and constrained to pass through $(0,0)$ and $(1,1)$.
+Empirical validation indicates a quartic function optimally captures this relationship:
 $f(x)=\alpha_{1}x^{4}+\alpha_{2}x^{3}+\alpha_{3}x^{2}+(1-\alpha_{1}-\alpha_{2}-\alpha_{3})x$ (3)
-To ensure reliable extrapolation, we calibrate the mapping curve using evaluation results of existing models as anchors.
-Our results show that the subset-to-full mapping generally maintains robustness across model architectures and training data, enabling the use of external models (e.g., Qwen2-72B [39]) as anchors for most tasks.
-We conduct corresponding experiments in Section 5.5.
-For data-sensitive tasks, models with similar training distributions provide more reliable anchors, indicating that data consistency takes precedence over architectural variation ensuring mapping accuracy.
-This calibration strategy enables accurate metric predictions for the complete evaluation set while maintaining computational efficiency.
-Finally, combining Eq. 1 and Eq. 2, we get our final metric prediction $p=f\circ y(C_{0})$, where $C_{0}$ is the estimated computation of training the target LLM.
+To ensure reliability, we calibrate $f$ using evaluation results from existing models as anchors.
+This subset-to-full mapping generally demonstrates robustness across diverse model architectures and training data, often permitting the use of external models (e.g., Qwen2-72B [40]) as anchors for many tasks (see Sec.
+5.3.3 for experiments).
+For data-sensitive tasks, however, models with similar training distributions provide more reliable anchors, prioritizing data consistency for mapping accuracy.
+The final metric prediction for a target LLM with estimated training computation $C_{0}$ is then $p=f\circ y(C_{0})$, combining the cluster-wise extrapolation $y(C_{0})$ from Eq.
+2 with the mapping $f$.
 5 Experiments
@@
 In our experimental setup, we train several smaller versions of the target model architecture for prediction.
-These models vary in size but share similar training procedures, with the training data scaled proportionally to their sizes.
-Downstream evaluation sets.
-We adopt the following widely-used benchmarks as our target downstream tasks: For evaluation, we adopt the following widely-used benchmarks, shown in Table 1.
-Evaluation sets cover popular downstream tasks of the language model, including math, logic, coding, reading comprehension, professional knowledge, etc.
-Table 1: Information of evaluation datasets used in the study.
-Dataset GSM8K MATH BBH TriviaQA MBPP AGIEval DROP MMLU-pro
-[7] [15] [32] [20] [3] [43] [9] [36]
-Domain Math Math Reasoning Knowledge Coding Comprehensive Reading Comprehensive
-#Questions 1,319 5,000 6,511 17,944 500 8,063 9,536 12,032
-#Shots in Prompt 8 4 3 5 3 5 3 5
-All models are evaluated in a few-shot in-context learning manner, where they need to generate final answer labels based on given demonstrations and test inputs.
-We aligned our evaluation setups with LLaMa3 [10].
-Model training.
-To establish performance predictions for large language models, we conduct systematic experiments with a suite of smaller-scale models across different parameter counts and training data volumes, while controlling for other training configurations such as learning rate, batch size, and additional hyperparameters.
-All models are trained on a constant learning rate scheduler and data with the same distribution.
-We list model configurations in Table 2.
-Table 2: Model architecture specifications across different sizes.
-122M 238M 411M 652M 973M 1.9B 7B 12B 70B (Target)
-Param. (M) 122 238 411 652 973 1,901 6,980 12,022 68,452
-Compute Per Token (B) 1.535 2.684 4.275 6.378 9.060 16.436 54.761 91.609 475.131
-Tokens (B) 26 45 72 108 153 277 923 1,544 8,012
-Layers 8 10 12 14 16 20 32 43 80
-Model Dimension 1,024 1,280 1,536 1,792 2,048 2,560 4,096 4,608 8,192
-FFN Dimension 3,584 4,480 5,376 6,272 7,168 8,960 14,336 16,128 28,672
-Heads 8 10 12 14 16 20 32 36 64
-KV Heads 8 10 12 14 16 20 8 12 8
+These models vary in size, but share similar training procedures, with the training data scaled proportionally to their sizes.
+We show the training configurations of the model in Appendix C.
+For evaluation, we adopt the following widely used benchmarks, including GSM8K [7], MATH [16], BBH [33], TriviaQA [21], MBPP [3], AGIEval [44], DROP [9], MMLU-pro [36].
+All models are evaluated in a few-shot in-context learning manner, and we aligned our evaluation setups with LLaMa3 [10].
+We evaluate our proposed COD performance scaling for LLMs against existing approaches on multiple public benchmarks.
+Using a series of smaller models with identical data distribution and architecture but varied configurations, we estimate the downstream task performance of a pretrained LLM with 70B parameters.
 5.2 Prediction Experiments
-Baselines.
-We evaluate our proposed COD performance scaling for LLMs against existing approaches.
-The evaluation is conducted on multiple public benchmarks mentioned above, where we utilize a series of smaller models with identical data distribution and architecture but different configurations to estimate the downstream tasks performance of the target 70B large language model.
-We compare against three existing prediction methods:
+We compare COD against three representative prediction methods:
   1. 1.
-    End-to-end performance-compute prediction: Extrapolate larger model metrics directly from smaller model evaluation set metrics using performance scaling laws.
+    End-to-end performance-compute prediction: Directly extrapolates large model metrics from smaller model evaluation set metrics using performance scaling laws.
   2. 2.
-    Passrate-compute prediction: Estimate large model passrates from smaller model passrates [1, 19].
-We repeat and evaluate 100 trials for each evaluation set to enhance the performance reliability on smaller models.
-For a fair comparison, we report the absolution prediction error on the passrate metric instead of greedy decoding accuracy.
+    Passrate-compute prediction [1, 20]: Estimates large model passrates from smaller model passrates.
+We conduct 100 trials per evaluation set for smaller models to enhance reliability and report absolute prediction error on the passrate metric.
   3. 3.
-    Loss-intermediate performance prediction: First predict the final training loss of large language model, then estimate downstream task metrics based on the relationship between smaller model evaluation metrics and their corresponding losses [6].
-We design two experimental groups to validate the benefits of clustering and the complete pipeline, respectively:
-  1. 1.
-    COD w/o.
-mapping: Performing difficulty-based clustering using K-Means, extrapolating within each cluster independently, and then aggregating metrics across clusters without requiring subset-to-full mappings.
-  2. 2.
-    COD complete: Complete multi-stage proposed approach consisting of clustering, predictable cluster filtering, subset extrapolation, and subset-to-full mapping.
-The comparative results across different benchmarks and estimation approaches are presented in Table 3.
-We evaluate prediction accuracy with the absolute error between predicted and actual performance.
-We report the prediction error on each single evaluation set and list the mean and the max prediction error.
-Table 3: Absolute prediction error on several evaluation sets.
-A prediction error less than 2 pp is considered an accurate estimate (marked in green), while an error greater than 5 pp is regarded as an invalid estimate (marked in red).
+    Loss-intermediate performance prediction [6]: First predicts the target LLM’s final training loss, then estimates downstream task metrics based on the relationship between smaller models’ evaluation metrics and their losses.
+Table 1: Absolute prediction error (in percentage points) on evaluation sets.
+Errors < 2% are considered accurate (green), while errors > 5% are considered invalid (red).
+$\downarrow$ indicates lower is better.
 Method Overall Metrics Individual Task Sets
@@
 COD (Complete) 1.63 2.38 2.23 1.28 1.77 1.64 2.19 2.38 0.23 1.35
-Results.
-Predictions with an absolute error of less than 2 percentage points (pp) are considered accurate estimations, and when the predicted values fall within the training metric fluctuation range, they are marked in green; predictions with an absolute error greater than 5 indicate invalid estimations, which reduce the overall reliability of the prediction method and are marked in red.
-These results show our approach significantly outperforms existing methods in both mean and maximum prediction errors, maintaining mean prediction error within 2 pp, thus offering practical guidance for large model training.
-While existing methods demonstrate good performance on certain evaluation sets, they consistently exhibit substantial estimation errors on a minority of sets, undermining the credibility of their predictions.
-Figure 4: Performance-compute relationship across difference prediction method
-Through visualization of the performance-compute relationship, we illustrate the distinctive characteristics of different prediction methods, shown in Fig.
-4.
-On the BBH evaluation set, while all three methods yield comparable estimates, end-to-end and loss-intermediate methods demonstrate inadequate fitting for small model evaluation points.
-In contrast, the cluster method reveals a more sophisticated and well-fitted multi-phase trajectory.
-For MATH and MMLU pro evaluation sets, the critical challenge involves determining whether large model metrics will experience accelerated growth with increased computing power or encounter performance plateaus.
-The loss-intermediate method exhibits an underestimation of the model capability ceiling, while the end-to-end method shows prediction errors exceeding 3 pp.
-The clustering method’s effectiveness can be attributed to its comprehensive analysis of evaluation set difficulty distributions and scaling laws.
-It successfully predicts growth patterns in mathematical evaluation sets where most problems demonstrate expanded improvement potential, while accurately capturing the diminishing scaling properties in evaluation sets with score saturation as computational resources increase.
-5.3 Comparison of Clustering Methods
-We evaluate the impact of clustering methods on the estimation approach.
-Our goal is to control the average distance between samples and cluster centers within clusters, making difficulty features more similar within clusters.
-We also ensure that the minimum number of questions in any cluster is not less than 10, considering that too small clusters may lead to instability in metric values.
-We compared our proposed Improved-MeanShift algorithm with clustering methods including DBScan, MeanShift, and K-Means.
-Since standard K-Means lacks the ability to filter outliers and directly control intra-cluster distances, we made the following adjustments: (1) Search for the number of clusters such that the minimum cluster size is close to but not less than 10 samples; (2) Draw spheres around cluster centers with a given threshold radius, and treat samples not covered by any sphere as outliers.
-If a cluster drops to less than 10 samples, we treat its samples as outliers.
-We use Intra-cluster Average Distance (IAD) and Outlier Rate (OR) as direct evaluation metrics.
-With similar OR, a smaller IAD indicates better clustering performance, as shown in Table 4.
-Additionally, we measure the benefits of different clustering methods on the prediction process by comparing the Extrapolation Errors(EE) of the predictable subset and Final prediction Errors (FE) after clustering, as shown in Table 5.
-Table 4: Clustering performance on popular benchmarks.
+We also evaluate two variants of our COD approach to validate the benefits of its components:
+  1. 1.
+    COD (w/o mapping): Employs difficulty-based clustering, extrapolates within each cluster, and aggregates metrics, bypassing the subset-to-full mapping stage.
+  2. 2.
+    COD (Complete): Our full proposed multi-stage approach, including clustering, predictable cluster filtering, subset extrapolation, and subset-to-full mapping.
+Comparative results are shown in Table 1.
+Prediction accuracy is measured by the absolute error between predicted and actual performance.
+We report mean and max prediction errors across all evaluation sets, as well as errors for individual sets.
+Our complete COD approach significantly outperforms existing methods in both mean (1.68%) and maximum (2.38%) prediction errors, offering reliable guidance for large model training.
+While baselines perform well on some sets, their substantial errors on others undermine overall credibility.
+Figure 4: Performance-compute relationship for different prediction methods on BBH, MATH, and MMLU-pro evaluation sets.
+Fig.
+4 visualizes the performance-compute relationships.
+On BBH, while all methods give similar final estimates, end-to-end and loss-intermediate methods fit small model data poorly.
+COD reveals a more complex, better-fitted multi-phase trajectory.
+On MATH and MMLU-pro, where predicting accelerated growth versus plateaus is crucial, the loss-intermediate method underestimates model ceilings, and the end-to-end method exceeds 3% error.
+COD’s superior performance stems from its nuanced analysis of difficulty distributions and scaling laws, allowing it to predict growth in challenging sets (e.g., MATH) and capture diminishing returns in saturated sets.
+The COD method is also capable of predicting LLMs’ performance after continual training phase on high-quality data.
+We conduct experiments and analysis on continual-trained LLMs and report prediction results in Appendix D.
+5.3 Ablation Study
+5.3.1 Comparison of Clustering Methods
+We assess the impact of different clustering algorithms on prediction accuracy.
+The goal is to achieve tight intra-cluster difficulty similarity (low average distance to center) while maintaining cluster stability (min.
+10 samples/cluster).
+We compare our Improved-MeanShift with DBScan, MeanShift, and K-Means.
+For K-Means, we adjust it to approximate our goals: (1) search for $k$ yielding min.
+cluster sizes $\approx 10$; (2) treat samples outside a radius (e.g., 2x average intra-cluster distance) from any cluster center as outliers, ensuring clusters don’t drop below 10 samples.
+We term this "Improved-KMeans" for this comparison.
+Clustering quality is measured by Intra-cluster Average Distance (IAD) and Outlier Rate (OR) (Table 2).
+Prediction benefits are measured by Extrapolation Error (EE) on the predictable subset and Final prediction Error (FE) (Table 3).
+[…]
 Method MMLU-pro GSM8k MATH BBH
@@
 Improved-MeanShift 0.2225 4.40 0.1854 4.93 0.1463 2.66 0.2143 33.58
-Table 5: Prediction errors across clustering algorithms.
+[…]
 Method Mean Max MMLU-pro GSM8k MATH BBH
@@
 Improved-MeanShift 1.23 1.66 2.20 2.23 1.27 1.35 2.20 2.23 1.14 1.28 0.29 1.77
-Table 4 shows that Improved-KMeans and Improved-MeanShift achieve better clustering performance, which is attributed to their incorporation of intra-cluster distance constraints during the clustering process.
-From Table 5, we can observe that the two methods with better clustering performance correspond to smaller extrapolation errors of estimable subsets and final metric prediction errors.
-Although Improved-KMeans achieves optimal clustering performance, its downstream estimation performance on GSM8k is notably inferior compared to other evaluation sets.
-We believe this is because the K-Means algorithm requires pre-specifying a more explicit number of clusters, but the number of clusters in evaluation sets is difficult to know in advance.
-While the number of clusters obtained through our search method is effective for some evaluation sets, it lacks stability and ultimately leads to excessive prediction errors in a few evaluation sets.
-In contrast, our adopted Improved-MeanShift algorithm inherently does not require pre-specifying the number of clusters; instead, it automatically determines this based on our intra-cluster distance constraints.
-This results in a more stable clustering performance and yields the smallest maximum estimation error across evaluation sets.
-We present additional clustering experimental results on more evaluation sets in Section C.1, where the conclusions are consistent with existing evaluation sets.
-5.4 Extrapolation Formula
-To evaluate the effectiveness of different fitting formulas, we conducted an ablation study comparing various formulations of the accuracy-compute relationship.
-Our baseline formula incorporates random guess probability, exponential decay, and a constant offset term:
-$f(C)=g+(1-g)*e^{-aC^{-b}-c}$ (3)
-To understand the contribution of each component, we perform ablation experiments by removing or modifying different terms.
-1) Without random guess component: $f_{1}(C)=e^{-aC^{-b}-c}$; 2) Without constant term $c$: $f_{2}(C)=g+(1-g)*e^{-aC^{-b}}$; 3) Direct power law relationship [19]: $f_{3}(C)=e^{-aC^{-b}}$.
-The comparative results of these formulations are shown in Table 6.
-Results from three evaluation sets, BBH, Math, and MMLU-pro, are presented here, showing the Extrapolation Error of extrapolatable clusters (EE), the Task Ratio of predictable subset (TR), and the Final prediction Error (FE).
-These results show that the proposed formula $f$ consistently achieves the smallest extrapolation error and final prediction error, while the ratio of estimable subsets remains similar across different clustering methods.
-Table 6: Ablation study results across different benchmarks.
+Table 2 shows Improved-KMeans and Improved-MeanShift yield better clustering (lower IAD) due to their intra-cluster distance constraints.
+Table 3 confirms these methods lead to lower EE and FE.
+Although Improved-KMeans has the best IAD, its GSM8k prediction is poor.
+This is likely because K-Means requires pre-specifying $k$, and our search for $k$ can be unstable, leading to large errors on some sets.
+In contrast, our Improved-MeanShift, which automatically determines $k$ based on distance constraints, offers more stable clustering and the lowest maximum prediction error.
+Additional results in Sec.
+E.1 support these conclusions.
+5.3.2 Different Extrapolation Formulas
+Table 4: Ablation study of extrapolation formulas. EE, TR, FE shown for BBH, MATH, MMLU-pro.
 Method BBH MATH MMLU-pro
 […]
-Direct Power Law 8.90 49.06 8.88 3.81 81.46 3.35 4.30 95.15 4.27
-w/o Random Guess 10.27 45.75 11.20 4.04 81.46 3.55 4.40 95.05 4.37
-w/o Constant c 2.14 57.26 4.01 1.40 81.46 1.56 3.85 95.60 3.88
-Ours 0.29 52.46 1.77 1.14 81.24 1.28 1.27 94.38 1.35
-In the control group, $f_{1}$ performs poorly on tasks with finite answer sets, where small models achieve non-zero scores that $f_{1}$ cannot effectively fit.
-$f_{2}$ removes $c$, which determines the maximum value of the prediction curve, assuming that the evaluation set performance would reach perfect scores given sufficient computation and parameters.
-However, this assumption is unreasonable due to the limited training data distribution and ambiguous answers in evaluation set questions, leading to inaccurate predictions.
-Direct power-law fitting $f_{3}$ fails to model both the metric range constraint of 0 to 1 and the characteristic that metric improvement is typically more difficult near Random Guess capability and capability saturation compared to other regions.
-We also observe that TR has little influence on prediction error, which also indicate the robustness of the proposed method.
-Some evaluation set with low TR due to the non-emergent subset, and our results show that this subset can largely be predicted with metrics of the extrapolatable clusters.
-We visualize the difficulty distribution of predictable subset and the full evaluation set in Appendix D.
-5.5 Anchor Point in Interpolation Mapping
-During the mapping phase from estimable subset metrics to full evaluation set metrics, we discovered that models with different training data and architectures exhibit similar mapping relationships.
-This allows us to leverage metrics from pre-trained models to refine the mapping relationship, thereby improving the accuracy of our final metric estimation.
-We use both Qwen2-72B [39] and an in-house model $M$ with a Mixture-of-Experts(MoE) [22] structure trained on the same data distribution as anchor points in the mapping phase.
-We first obtain the interpolation curve using only small model metrics with points (0,0) and (1,1), then verify the compatibility of anchor points with the existing interpolation curve.
-When the score of the full set is 0 (1), the score of the subset must also be 0 (1).
-Results show that despite differences in computational requirements, model architectures, and training data, these models share similar mapping relationships.
-This finding also indicates that estimable subset metrics are highly correlated with full-set metrics.
-Compared to loss-intermediate estimation, estimable subset metrics maintain predictability while reducing interference from other model parameters.
-Based on these observations, we incorporated the mapping relationships from pre-trained models into the interpolation process, thereby improving both the accuracy and confidence of our large model metric estimations.
-We establish three experimental configurations of our method:
-  • COD w/o. anchor: The complete estimation process is employed except for not using anchor point interpolation in the Mapping phase.
-  • COD w.
-out-of-distribution (OOD) anchor: The complete proposed methodology incorporates both difficulty-based clustering and predictable subset identification.
-Using the 72B Qwen2 pretraining model as the anchor model [39].
-  • COD w. in-domain(ID) anchor: Using an in-house MoE model with consistent training distribution but different model architecture as the anchor point.
-We present the results in Table 7, demonstrating that incorporating both out-of-distribution models and in-distribution models as anchors consistently enhance prediction accuracy.
-These findings suggest a relatively stable correlation between the metrics of predictable subsets and the full dataset, indicating that the relationship between subset and full-set metrics remains consistent across models trained on different data and with varying architectures.
-This property enables us to leverage evaluation results from existing models to improve the accuracy of metric predictions for large models with new data and structures.
-Besides, the division of the evaluation set obtained through clustering is an intrinsic property of the evaluation set itself, independent of the model architecture and training data.
-Therefore, the predictable subset derived from clustering can also be extended to estimate the metrics of new models.
-Additionally, we conduct the ablation study on the interpolation method in Section C.2, and results indicate that quartic functions are suitable in our setting.
-Table 7: Influence of anchor point usage in the mapping stage.
+Direct Power Law ($f_{3}$) 8.90 49.06 8.88 3.81 81.46 3.35 4.30 95.15 4.27
+w/o Random Guess ($f_{1}$) 10.27 45.75 11.20 4.04 81.46 3.55 4.40 95.05 4.37
+w/o Constant c ($f_{2}$) 2.14 57.26 4.01 1.40 81.46 1.56 3.85 95.60 3.88
+Ours ($f$) 0.29 52.46 1.77 1.14 81.24 1.28 1.27 94.38 1.35
+We ablate our proposed fitting formula $f(C)=g+(1-g)\cdot e^{-aC^{-b}-c}$ (Ours) by removing or modifying components: 1) $f_{1}(C)=e^{-aC^{-b}-c}$ (w/o random guess); 2) $f_{2}(C)=g+(1-g)\cdot e^{-aC^{-b}}$ (w/o constant $c$); 3) $f_{3}(C)=e^{-aC^{-b}}$ (Direct Power Law [20]).
+Table 4 shows Extrapolation Error (EE), Task Ratio of predictable subset (TR), and Final prediction Error (FE).
+Our proposed formula $f$ consistently achieves the lowest EE and FE.
+$f_{1}$ struggles with finite-answer tasks where small models have non-zero scores.
+$f_{2}$ inaccurately assumes perfect scores are attainable, ignoring data limitations and task ambiguities.
+The direct power law ($f_{3}$) fails to model the 0-1 metric range and the varying difficulty of improvement near random guess and saturation.
+TR’s minor influence on prediction error suggests robustness; even with low TR due to non-emergent tasks, their performance can be largely predicted via mapping (difficulty distributions visualized in Appendix F).
+5.3.3 Different Anchor Points in Interpolation Mapping
+We find that the mapping relationship from predictable subset metrics to full evaluation set metrics is similar across models with different training data and architectures.
+This allows leveraging pre-trained models as "anchors" to refine the mapping and improve final estimation accuracy.
+We use Qwen2-72B [40] (OOD anchor) and an in-house MoE model [23] with the same training data but different architecture (ID anchor).
+We first derive an interpolation curve using only small model metrics and fixed points $(0,0)$, $(1,1)$, then assess anchor compatibility.
+This shared mapping implies estimable subset metrics are highly correlated with full-set metrics and less prone to interference from other model parameters than loss-intermediate predictions.
+We test three configurations:
+  • COD (w/o anchor): Full COD pipeline, but no anchor points in the mapping phase;
+  • COD (w. OOD anchor): COD with Qwen2-72B as an out-of-distribution anchor;
+  • COD (w. ID anchor): COD with our in-house MoE model as an in-domain anchor.
+Table 5: Influence of anchor points in the mapping stage on prediction error (pp). Lower is better.
 Method Overall Individual Task Sets
-Mean Max GSM8k MATH BBH TriviaQA MBPP AGIEval DROP MMLU-pro
+[…]
 w/o. anchor 3.96 8.80 2.17 5.46 5.08 1.68 8.80 2.38 4.44 1.68
@@
 w. ID anchor 1.63 2.38 2.23 1.28 1.77 1.64 2.19 2.38 0.23 1.35
+Table 5 shows that incorporating either OOD or ID anchors consistently enhances prediction accuracy.
+This suggests a stable correlation between predictable subset and full-set metrics across diverse models, enabling the use of existing model evaluations to improve predictions for new models.
+Furthermore, since our clustering identifies intrinsic properties of evaluation sets, the derived predictable subsets are applicable to new models.
+Additional ablations on the interpolation method (Sec.
+E.2) confirms quartic functions are suitable.
 6 Conclusion and Discussion
-In this work, we introduce a novel downstream performance scaling framework including (1) a difficulty-based clustering approach that effectively models the underlying distribution of each evaluation set; (2) a theoretically grounded scaling law for downstream task performance that provides a fitting formula for performance-compute extrapolation; and (3) a systematic methodology for identifying and leveraging predictable subset that provides a robust intermediate metric for accurate full-set performance predictions.
-Our framework, while effective for dense transformers, has not been fully explored for cost-efficient MoE models and does not account for the annealing phase in training, where high-quality data can rapidly enhance performance.
-The COD method requires sufficient test cases and is not suited for multiple-choice tasks, where performance metrics may diverge from true passrates.
-Additionally, the framework’s theoretical foundation is insufficient for chain-of-thought reasoning, necessitating future adaptations to address these challenges.
-We provide detailed discussion in Appendix E.
+In this work, we introduce a novel downstream performance scaling framework including (1) the COD method that effectively models the underlying diverse scaling patterns of evaluation tasks; (2) a scaling law for downstream task performance that provides a fitting formula for performance-compute extrapolation; and (3) a systematic methodology for identifying and leveraging predictable subset that provides a robust intermediate metric for accurate full-set performance predictions.
+We discuss the limitations of the COD method in Appendix G.
 Looking forward, our approach can be further expanded across model architectures, training methods, and evaluation set types, while extending this framework to address chain-of-thought reasoning patterns offer promising avenues for future research.
+Boarder Impact
+The LLMs performance scaling prediction method proposed in this work can be used to guide decision-making for training LLMs and to monitor whether the LLM training process aligns with expectations.
+Accurate prediction allows for better planning of compute and human resources, thereby playing a valuable role in scientifically advancing the training of larger LLMs.
+However, if the prediction results are inaccurate, it may lead to outcomes that do not meet expectations for the invested resources, potentially resulting in resource waste or economic losses.
 Appendix A Improvements of Clustering Algorithm
@@
 6:         Calculate distance $dist_{i}$ to its cluster center
-7:         if $dist_{i}>R$ then
+7:         if $\mathrm{dist}_{i}>R$ then
 8:              Reset label to $-1$
@@
 17:until no label changes
-Filtering Zero-performance Samples
+Filtering zero-performance samples.
 In the evaluation set, there may exist a few extremely difficult problems that require sufficient model parameters to emerge.
@@
 Therefore, we pre-filter these zero-performance samples before clustering, treating them as outliers that do not participate in the clustering process.
-This approach eliminates the need to consider their metrics under large models during subsequent extrapolation and prevents disruption to the clustering of normal difficult samples.
+This approach obviates the necessity of considering their metrics under large models during subsequent extrapolation, and prevents disruption to the clustering of the remaining samples.
 A.2 Smoothing Techniques
-Horizontal smoothing: adjacent checkpoint smoothing.
 Metric fluctuations of individual samples in downstream tasks are not solely due to limited sampling.
@@
 This approach also reduces the number of zero-performance samples, further improving clustering and prediction effectiveness.
-Vertical smoothing.
-Each sample’s features represent the expected correct response rate across models of increasing size, forming a partially ordered sequence.
-However, the Euclidean distance used for measurement does not consider this sequential information.
-For example, if a cluster center has a feature sequence of $[0,0,0,0.5]$, sample A with $[0,0,0.2,0.5]$ and sample B with $[0.2,0,0,0.5]$, sample A clearly fits the cluster better than sample B, yet their Euclidean distances are identical.
-Note that this smoothing method may not be effective for all downstream tasks.
-Our current observations suggest that for freeform tasks with limited solution spaces (such as multiple choice, ordering, or judgment questions in freeform format), once models learn to answer within the solution space, their random guess metrics on the evaluation set will be non-zero, more significantly affected by recent training batch data and few-shot cases in prompts.
-In such cases, vertical smoothing is more likely to bring positive benefits.
-In our experiment, we only adopt horizontal smoothing, and leave vertical smoothing as an optional selection.
-Appendix B Proof of Proposition
-We use Proposition B.1 to derive scaling law for downstream task performance (Proposition B.2).
-Proposition B.1 (Arithmetic-geometric mean difference).
+Appendix B Proof of Theorem 1
+We use lemma 1 to derive the scaling law for downstream task performance (theorem A1).
+Lemma 1 (Arithmetic-geometric mean difference).
 For any sequence of positive real numbers $\{x_{i}\}_{i=1}^{n}$, let:
@@
 Then the difference between the arithmetic mean and geometric mean can be estimated as:
-$\Delta=\mu_{a}-\mu_{g}=\frac{1}{n}\sum_{i=1}^{n}x_{i}-\left(\prod_{i=1}^{n}x_{i}\right)^{\frac{1}{n}}=\frac{\sigma^{2}}{2\mu_{a}}+o(\mu_{a})$ (4)
+$\Delta$ $=\mu_{a}-\mu_{g}=\frac{1}{n}\sum_{i=1}^{n}x_{i}-\left(\prod_{i=1}^{n}x_{i}\right)^{\frac{1}{n}}=\frac{\sigma^{2}}{2\mu_{a}}+o(\mu_{a}).$ (4)
 Proof.
@@
 Using Taylor expansion of $\log x$ around $\mu$:
-$\log x=\log\mu+\frac{x-\mu}{\mu}-\frac{(x-\mu)^{2}}{2\mu^{2}}+o\left((x-\mu)^{2}\right)$ (6)
+$\log x=\log\mu+\frac{x-\mu}{\mu}-\frac{(x-\mu)^{2}}{2\mu^{2}}+o\left((x-\mu)^{2}\right).$ (6)
 We can simplify:
-$\log(GM)$ $=\frac{1}{n}\sum_{i=1}^{n}\log x_{i}$
-$=\log\mu+\frac{1}{n}\sum_{i=1}^{n}\left(\frac{(x_{i}-\mu_{a})}{\mu_{a}}-\frac{(x_{i}-\mu_{a})^{2}}{2\mu_{a}^{2}}\right)+o(\mu_{a})$
+$\frac{1}{n}\sum_{i=1}^{n}\log x_{i}=$ $\log\mu+\frac{1}{n}\sum_{i=1}^{n}\left(\frac{(x_{i}-\mu_{a})}{\mu_{a}}-\frac{(x_{i}-\mu_{a})^{2}}{2\mu_{a}^{2}}\right)+o(\mu_{a})$
 $=$ $\log\mu+\frac{1}{\mu}\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}-\mu_{a}\right)}_{\text{equal to 0}}+\frac{1}{2\mu_{a}^{2}}\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu_{a})^{2}\right)}_{\sigma^{2}}+o(\mu_{a})$
-$=\log\mu-\frac{\sigma^{2}}{2\mu^{2}}+o(\mu_{a})$
+$=$ $\log\mu-\frac{\sigma^{2}}{2\mu^{2}}+o(\mu_{a}).$
 Therefore:
-$\mu_{a}-\mu_{g}=\mu_{a}\left(1-\exp{\left(-\frac{\sigma^{2}}{2\mu^{2}}\right)}\right)+o(\mu_{a})$ (7)
+$\mu_{a}-\mu_{g}=\mu_{a}\left(1-\exp{\left(-\frac{\sigma^{2}}{2\mu^{2}}\right)}\right)+o(\mu_{a}).$ (7)
 When $\frac{\sigma^{2}}{2\mu^{2}}$ is small, this can be approximated as:
 $\Delta\approx\frac{\sigma^{2}}{2\mu_{a}}$ (8)
-Proposition B.2 (Scaling law for downstream task performance).
+Theorem A1 (Scaling Law for Downstream Task Performance).
 Given a language model trained with computational budget $C$, and a set of downstream tasks $P$, under the following assumptions:
   1. 1.
-    The relationship between the answer loss and compute follows a power law, $\frac{1}{n}\sum_{(q,\mathrm{ans})}\mathrm{loss}_{\mathrm{ans}}(C)\sim aC^{-b}+c$;
+    Given a task $(q,\mathrm{ans})\in P$, the relationship between the answer loss and the compute follows a power law, $\frac{1}{n}\sum_{(q,\mathrm{ans})}\mathrm{loss}_{\mathrm{ans}}(C)\sim aC^{-b}+c$;
   2. 2.
-    For tasks with a finite answer set, the model gives a random guess choice if it cannot truly solve it;
+    For tasks with a finite answer set, the model gives a random guess choice if it cannot truly solve it, i.e. $\mathbb{E}[\mathrm{Acc|incorrect}]=g$;
   3. 3.
-    Task passrate equals the product of the predicted probability of each token $p_{\mathrm{ans}}=\prod_{t\in\mathrm{ans}}p(t)$, which means that each task has unique answer and model output answer only without thinking progress.
+    Task passrate equals the product of the predicted probability of each token $p_{\mathrm{ans}}=\prod_{t\in\mathrm{ans}}p(t)$, which means that each task has a unique answer and model output answer only without thinking progress.
 The expected accuracy on tasks $P$ can be modeled as:
-$\mathbb{E}_{p}[\mathrm{Acc}(C)]=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{2\mu}\right)+o(\mu)$ (9)
-where:
-  • $g$ represents the random guess performance floor;
-  • $a,b,c$ are positive constants;
-  • $\mu=\frac{1}{\#P}\sum_{(q,\mathrm{ans})\in P}loss_{\mathrm{ans}}$;
-  • $\sigma^{2}=\frac{1}{\#P}\sum(\mathrm{loss}_{\mathrm{ans}\_t}-\mu)^{2}$.
+$\mathbb{E}_{p}[\mathrm{Acc}(C)]=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{\mu}\right)+o(\mu),$ (9)
+where $g$ represents the random guess performance floor, $1-g$ represents the maximum achievable performance improvement, $a,b,c$ are positive constants, and $\mu=\frac{1}{\#P}\sum_{(q,\mathrm{ans})\in P}\mathrm{loss}_{\mathrm{ans}}$; $\sigma^{2}=\frac{1}{\#P}\sum(\mathrm{loss}_{\mathrm{ans}_{t}}-\mu)^{2}$.
 Proof.
-We first use assumption 3 to establish the relationship between model passrate and loss on a task.
-$-\log(p_{\mathrm{ans}})=-\log\left(\prod_{t\in\mathrm{ans}}p(t)\right)=-\sum_{t\in\mathrm{ans}}\log(p_{t})=\mathrm{loss}_{\mathrm{ans}}$ (10)
+We first use Assumption 3 to establish the relationship between model passrate and loss on a task:
+$-\log(p_{\mathrm{ans}})$ $=-\log\left(\prod_{t\in\mathrm{ans}}p(t)\right)=-\sum_{t\in\mathrm{ans}}\log(p_{t})=\mathrm{loss}_{\mathrm{ans}}.$ (10)
 Then take the exponential of both sides, and then take the expectation with respect to different tasks in the evaluation set $p=(q,\mathrm{ans})\in P$.
 We note that both $p_{\mathrm{ans}}$ and $\mathrm{loss}_{\mathrm{ans}}$ are functions of $C$.
-$\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]$ $=\mathbb{E}_{p}[\exp(-\mathrm{loss}_{\mathrm{ans}}(C))]$ (11)
-$=\frac{1}{n}\sum_{(q,\mathrm{ans})\in P}\exp(-\mathrm{loss}_{\mathrm{ans}}(C)).$ (12)
-We can adopt Proposition B.1 to switch from arithmetic mean to geometric mean of $\mathrm{loss}$, and apply the power law assumption 1.
+$\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]$ $=\mathbb{E}_{p}[\exp(-\mathrm{loss}_{\mathrm{ans}}(C))]=\frac{1}{n}\sum_{(q,\mathrm{ans})\in P}\exp(-\mathrm{loss}_{\mathrm{ans}}(C)).$ (11)
+We can adopt lemma 1 to switch from arithmetic mean to geometric mean of $\mathrm{loss}$, and apply the power law assumption 1.
 $\frac{1}{n}\sum_{(q,\mathrm{ans}_{t})\in P}\exp(-\mathrm{loss}_{\mathrm{ans}}(C))=$ $\exp{\underbrace{\left(-\frac{1}{n}\sum_{(q,\mathrm{ans}_{t})\in P}\mathrm{loss}_{\mathrm{ans}}(C)\right)}_{\text{use loss scaling law}}}+\frac{\sigma^{2}}{2\mu}+o(\mu)$ (12)
@@
 Take these results into Eq.
-15, and we have:
+14, and we have:
 $\mathbb{E}_{p}[\mathrm{Acc}(C)]$ $=\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]+(1-\mathbb{E}_{p}[p_{\mathrm{ans}}(C)])*g$ (15)
@@
 $=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{2\mu}\right)+o(\mu)$ (17)
-Proposition B.2 demonstrates that a metric of an evaluation set with similar difficulty features can be effectively modeled using the following formula:
-$f(C)=g+(1-g)\exp{(-aC^{-b}-c)}$ (19)
-Appendix C Additional Ablation Studies
-C.1 Comparison of Clustering Methods on Extra Evaluation Sets.
-We provide additional clustering evaluation results across more evaluation sets in Table C1 and Table C2, which maintain consistency with the conclusions presented in the main text.
-Table C1: Clustering performance on advanced task benchmarks (IAD: Intra-cluster Average Distance, OR: Outlier Rate)
+theorem A1 demonstrates that a metric of an evaluation set with similar difficulty features can be effectively modeled using the following formula:
+$f(C)=g+(1-g)\exp{(-aC^{-b}-c)}.$ (18)
+Appendix C Experiment Settings
+Model training.
+To establish performance predictions for large language models, we conduct systematic experiments with a suite of smaller-scale models across different parameter counts and training data volumes, while controlling for other training configurations such as learning rate, batch size, and additional hyperparameters.
+All models are trained on a constant learning rate scheduler with a few-step warmup stage and data with the same distribution.
+We list model configurations in Table A1.
+Table A1: Model architecture specifications across different sizes.
+122M 238M 411M 652M 973M 1.9B 7B 12B 70B (Target)
+Param. (M) 122 238 411 652 973 1,901 6,980 12,022 68,452
+Compute Per Token (B) 1.535 2.684 4.275 6.378 9.060 16.436 54.761 91.609 475.131
+Tokens (B) 26 45 72 108 153 277 923 1,544 8,012
+Continue-Trained Tokens (B) 3 5 8 12 18 33 114 191 1,000
+Model Dimension 1,024 1,280 1,536 1,792 2,048 2,560 4,096 4,608 8,192
+FFN Dimension 3,584 4,480 5,376 6,272 7,168 8,960 14,336 16,128 28,672
+Heads 8 10 12 14 16 20 32 36 64
+KV Heads 8 10 12 14 16 20 8 12 8
+Evaluation settings.
+We conducted performance scaling estimation experiments across 8 major LLM evaluation sets.
+These evaluation sets span a diverse range of capabilities, including Math, Reasoning, Knowledge, Coding, Reading, and general abilities.
+All pretrained LLMs were evaluated using a few-shot methodology to obtain the performance metrics.
+Detailed information is provided in Table A2.
+Table A2: Information of evaluation datasets.
+Dataset Domain #Questions #Shots
+GSM8K Math 1,319 8
+MATH Math 5,000 4
+BBH Reasoning 6,511 3
+TriviaQA Knowledge 17,944 5
+MBPP Coding 500 3
+AGIEval Comprehensive 8,063 5
+DROP Reading 9,536 3
+MMLU-pro Comprehensive 12,032 5
+Appendix D Performance Prediction for Continue-Pretrained LLMs
+Leading industry pre-trained LLMs (e.g., Deepseek-v3 [25], Llama3 [14], Qwen-2.5 [39]) adopt the Continual Pre-Training (CPT) strategy of concentrating high-quality data towards the end of the pre-training process.
+This phase is typically accompanied by learning rate decay, enabling the model to fully absorb this high-quality data.
+Due to significant changes in data distribution and the learning rate schedule, this approach often yields substantial improvements in metrics.
+Predicting a large model’s final capability based solely on its performance during a “stable” phase with consistent data distribution does not reflect its ultimate capability.
+Therefore, we supplement this by providing metric predictions for the high-quality CPT phase.
+The relationship between model parameter scale and the volume of CPT tokens is listed in Table A1.
+We conduct the same COD pipeline for CPT models.
+We control the data distribution of the stable and decay phases for various smaller models, as well as their token-to-parameter ratio, to be consistent with the large model targeted for prediction.
+The last checkpoint is used for evaluation.
+Based on prior clustering labels, we perform fitting, extrapolation, and mapping to obtain the predicted performance for the large model.
+Table A3: Predicted vs. actual metrics for an LLM with 70B parameters after high-quality continued pretraining.
+Evaluation Set Predicted Metric Actual Metric Prediction Error
+MMLU-pro 61.58 59.34 2.24
+BBH 84.66 85.32 0.66
+TriviaQA 79.64 84.05 4.41
+MBPP 73.00 73.20 0.20
+AGIEval 61.66 64.18 2.52
+DROP 81.65 81.39 0.26
+gsm8k 86.88 91.81 4.93
+MATH 49.29 52.68 3.39
+Results listed in Table A3 show that the proposed COD method achieves an average prediction error of 2.33%.
+We observe that MATH, GSM8k, and TriviaQA exhibit relatively large prediction errors. We hypothesize that there are two main categories of reasons for this inaccuracy:
+  1. 1.
+    The CPT data and the evaluation sets possess a significant correlation.
+For example, in math-related evaluation sets, a modest amount of training can yield substantial improvements in performance metrics.
+In such scenarios, the metrics for smaller models tend to show greater volatility and have inaccurate evaluations.
+  2. 2.
+    The CPT data exhibits inherent distribution bias, such that certain evaluation sets, such as TriviaQA, do not derive performance gains from it.
+This leads to potential significant fluctuations in the metrics after the CPT phase, thereby diminishing the accuracy of extrapolating to larger models.
+Appendix E Additional Ablation Studies
+E.1 Comparison of Clustering Methods on Extra Evaluation Sets.
+We provide additional clustering evaluation results across the remaining evaluation sets in Table A4 and Table A5, which maintain consistency with the conclusions presented in the main text.
+Table A4: Clustering performance on advanced task benchmarks (IAD: Intra-cluster Average Distance, OR: Outlier Rate)
 Method TriviaQA AGIEval DROP MBPP
@@
 Improved-MeanShift 0.1871 11.54 0.2100 11.50 0.1974 19.88 0.1745 21.60
-Table C2: Prediction errors on advanced task benchmarks.
+Table A5: Prediction errors on advanced task benchmarks.
 Method Mean Max TriviaQA AGIEval DROP MBPP
@@
 Improved-MeanShift 1.13 1.61 1.58 2.38 1.58 1.64 1.11 2.38 0.26 0.23 1.56 2.19
-C.2 Interpolation Method
+E.2 Interpolation Method
 To evaluate different interpolation methods for prediction accuracy, we compared various mathematical approaches.
 Our baseline method uses quartic polynomial interpolation, which we compare against several alternative approaches, including Cubic spline interpolation, Cubic polynomial interpolation, and Quintic polynomial interpolation
-The comparative results across different benchmarks are shown in Table C3. We report the prediction error between the real performance of a large model and the mapping result.
-Table C3: Comparison of different interpolation methods across benchmarks.
+The comparative results across different benchmarks are shown in Table A6. We report the prediction error between the real performance of a large model and the mapping result.
+Table A6: Comparison of different interpolation methods across benchmarks.
 Prediction Error BBH Math MMLU-pro
@@
 Furthermore, in Fig.
-C1, we plot the mapping process using different interpolation formulas, where the x-axis represents the predictable subset indices and the y-axis represents the full set indices.
-The red points are the numerical values to be fitted, green points represent predicted values, purple points represent anchor points and blue points show the actual performance.
+A1, we plot the mapping process using different interpolation formulas, where the x-axis represents the predictable subset indices and the y-axis represents the full set indices.
+The red points are the numerical values to be fitted, green points represent predicted values, purple points represent anchor points, and blue points show the actual performance.
 The prediction performance shows certain robustness across different interpolation methods.
@@
 Based on the above results, we observe that the Cubic Polynomial shows larger prediction errors due to underfitting on BBH, while the Cubic Spline exhibits some overfitting.
-Both Quartic Polynomial and Quintic Polynomial perform well, therefore we chose the Quartic Polynomial method as it requires fewer fitting parameters.
-Figure C1: Performance mapping with different interpolation methods on BBH evaluation set.
+Both Quartic Polynomial and Quintic Polynomial perform well, therefore, we chose the Quartic Polynomial method as it requires fewer fitting parameters.
+Figure A1: Performance mapping with different interpolation methods on the BBH evaluation set.
 The cubic spline is overfitted, and the cubic polynomial method is underfitted.
-Quartic polynomials and quintic polynomials are comparable while quartic polynomial has fewer parameters.
-Appendix D Difficulty Distribution of Predictable Subset
+Quartic polynomials and quintic polynomials are comparable, while a quartic polynomial has fewer parameters.
+Appendix F Difficulty Distribution of Predictable Subset
 We analyze the proportion of predictable subset tasks across different difficulty levels.
 The difficulty distributions of predictable subset versus complete sets for different evaluation benchmarks are illustrated in Fig.
-D2.
+A2.
 We use the scores from the 12B model as the basis for difficulty classification.
-The results show that MMLU-pro and GSM8k evaluation sets have larger proportions of predictable subset, indicating that most questions in these datasets exhibit good performance scaling properties.
+The results show that MMLU-pro and GSM8k evaluation sets have larger proportions of predictable subsets, indicating that most questions in these datasets exhibit good performance scaling properties.
 In contrast, many difficult questions with near-zero scores in the MATH evaluation set fall outside the predictable subset, requiring adjustment during the mapping phase.
-Meanwhile, BBH shows consistent proportions of predictable subset across difficulty levels, as some of its questions demonstrate oscillatory patterns with limited improvement despite increased computing.
-The proportion of predictable subset can serve as a metric for assessing evaluation set quality.
-Evaluation sets with larger predictable subset yield more reliable experimental conclusions from smaller models.
+Meanwhile, BBH exhibits consistent proportions of its predictable subset across varying difficulty levels, as some questions display oscillatory patterns with limited improvement, even with increased computational resources.
+The proportion of the predictable subset can serve as a metric for assessing evaluation set quality.
+Evaluation sets with larger predictable subsets yield more reliable experimental conclusions from smaller models.
 When constructing evaluation sets, we recommend screening or supplementing unpredictable clusters and ensuring a minimum number of questions for each difficulty feature to reduce metric volatility.
-Figure D2: Difficulty distribution comparison on a 12B model between predictable subset and full evaluation set.
-Appendix E Limitations
-Influence of model structure and training configurations.
+Figure A2: Difficulty distribution comparison on a 12B model between predictable subset and full evaluation set.
+Appendix G Limitations
+Influence of model structure.
 Mixture-of-Experts (MoE) models excel in training and inference cost, and are widely used in production.
@@
 However, we believe that the proposed method is not significantly affected by the model architecture.
-If we apply the complete pipeline to MoE models, we expect to achieve similar results.
-Besides, we only study the pre-training performance prediction with a constant learning rate and do not cover the impact of the annealing training.
-In this phase, higher-quality data is usually adopted, which can rapidly improve the model’s capabilities.
-As a result, performance prediction faces greater challenges.
+We expect similar results for MoE models.
 Category of evaluation sets.
 The proposed Clustering-on-Difficulty method requires a sufficient number of test cases, as too few samples can lead to unstable cluster metrics and ineffective estimation.
-However, from an evaluation set design perspective, an evaluation set with good predictive properties enables more effective generalization from small-scale to large-scale models, thus providing better guidance for model iteration.
-Furthermore, for multiple-choice tasks, the model only needs to assign a higher probability to the correct option compared to others, creating a discrepancy between this metric and the model’s true passrate.
-Given that more evaluation sets are adopting the Chain-of-Thoughts prompts, we have not included multiple-choice tasks that only require option selection.
+From an evaluation set design perspective, an evaluation set with good predictive properties enables more effective generalization from small-scale to large-scale models, thus providing better guidance for model iteration.
+Furthermore, for multiple-choice tasks, the model only needs to assign a higher probability to the correct option compared to others, creating a discrepancy between the answer loss and the model’s true passrate.
+Given that more evaluation sets are adopting Chain-of-Thoughts prompts, we have not included multiple-choice tasks that only require option selection.
 Chain-of-thought performance prediction.
-Proposition B.2 assumes that evaluation sets directly assess models’ ability to provide answers.
+theorem A1 assumes that evaluation sets directly assess models’ ability to provide answers.
 However, increasingly more evaluations allow models to think before providing answers.

### v3

@@
 […]
-Chenyin Xu at , Ke Shen at
+Chengyin Xu at , Ke Shen at
 1 Introduction
@@
 One line of work attempts to extrapolate the performance of a large model by modeling the performance-loss relationship [8, 16, 11, 43, 31], but this often fails to capture the emergent behaviors of LLMs and the mismatch between the in-domain loss and downstream metrics [49].
-Another line of research focuses on direct extrapolation of the performance-compute relationship [1, 20], yet the uneven difficulty distribution between different evaluation samples undermines its accuracy.
+Another line of research focuses on direct extrapolation of the performance-compute relationship [1, 23], yet a single family of curves usually fails to capture the performance on evaluation benchmarks with complex difficulty distributions across samples.
 A critical oversight in current approaches is the assumption of uniform scaling behavior across diverse evaluation samples.
 We observe that different evaluation samples actually follow distinct performance scaling patterns, and thus applying a single extrapolation formula to the entire evaluation set is suboptimal.
-We give a detailed analysis in Sec.
-3.
+We give a detailed analysis in Sec.˜3.
 To address these challenges, we propose a new performance scaling law, derived from the existing loss scaling law [25], specifically applicable to evaluation subsets that exhibit consistent performance scaling patterns.
@@
 We validate our COD approach on eight popular evaluation sets, including MATH [19], BBH [38], and MMLU pro [41] datasets.
-COD achieves an average prediction error of 1.36% on an LLM with 70B parameters.
+COD achieves an average prediction error of 1.36%ha on an LLM with 70B parameters.
 Our results demonstrate that this difficulty-aware framework substantially outperforms existing methods, establishing a promising paradigm for accurate downstream performance scaling of LLMs.
@@
 2.2 Downstream Task Performance Scaling
-Predicting downstream task performance remains a critical challenge due to emergent abilities in LLMs that manifest only after exceeding task-specific thresholds [37, 29].
+Predicting downstream task performance remains a critical challenge due to emergent abilities in LLMs that some capabilities manifest only after exceeding task-specific thresholds [42, 32].
 Recent works, such as using loss as a proxy [8] or increasing metric resolution [23], have demonstrated potential but encounter challenges in aligning surrogate metrics with original task objectives.
@@
 End-to-end performance-compute prediction.
-These methods [20, 28, 1] directly model the relationship between performance and compute (or the number of model parameters).
-Additionally, Achiam et al.
+These methods [23, 31, 1, 7] directly model the relationship between performance and the compute budget (or the number of model parameters).
+Specifically, Achiam et al.
 [1] estimate and fit this relationship using a subset of the evaluation set, while still failing to predict the full set.
@@
 [23] address the challenge of non-emergent capabilities in smaller models by employing multiple non-greedy decoding evaluations, thereby enabling accurate extrapolation of performance predictions for models with up to 2.4B parameters.
+Caballero et al.
+[7] propose a smooth broken power-law that models LLM scaling by decomposing it into multi-segment power laws.
+However, when predicting metrics for large-scale models (e.g., 70B parameters), performance trends often exhibit unexpected inflection points due to emergent capabilities or saturation effects, making piecewise functions inadequate for capturing these novel scaling regimes.
 3 Pilot Study
@@
 Moreover, training configurations, such as model size and learning rate, can significantly affect not only the final loss but also the model’s generalization capabilities.
-Fig.
-1(left) illustrates the performance–loss relationships for LLMs of different sizes on the CEval benchmark [30].
+figure˜1(left) illustrates the performance–loss relationships for LLMs of different sizes on the CEval benchmark [33].
 At the same training loss level, smaller models can outperform larger ones in terms of test accuracy.
 Because smaller models initially exhibit weaker in-domain fitting capacity, they typically require more training steps to reach the same loss value, which can lead to better in-domain generalization once they do.
-Fig.
-1(middle) compares the performance of LLMs trained under different learning rate schedules on the GSM8k dataset [7].
+figure˜1(middle) compares the performance of LLMs trained under different learning rate schedules on the GSM8k dataset [9].
 At the same loss level, the performance under the cosine schedule is always worse than that under the constant schedule, indicating that a lower learning rate may prioritize memorization over generalization, thereby diminishing downstream performance.
@@
 Different task samples exhibit unique computational thresholds, learning slopes, and upper bounds, making it challenging to find a single fitting function (or function group) that generalizes well across diverse task samples.
-Fig.
-1(right) illustrates the performance-compute relationships on three random clusters of the BBH benchmark [33], with each cluster containing samples with similar difficulty.
+figure˜1(right) illustrates the performance-compute relationships on three random clusters of the BBH benchmark [38], with each cluster containing samples with similar difficulty.
 Even within a single evaluation set, these scaling curves can vary significantly, indicating that a one-size-fits-all performance-compute curve is insufficient for capturing the full spectrum of a downstream evaluation set.
@@
 4 Method
-In this section, we introduce the COD method in four parts, as illustrated in Fig.
-2: 1) We show the advantages of COD and present an improved mean-shift clustering algorithm (Sec.
-4.1); 2) We derive a performance scaling law corresponding to task difficulty variance, which supports extrapolating the performance-compute relationship for task clusters with similar difficulty features (Sec.
-4.2).
-We fit cluster-wise performance-compute curves on small models and filter extrapolatable clusters; 3) We extrapolate the performance on extrapolatable clusters and predict the accuracy of the target large model on the predictable subset(Sec.
-4.3); 4) We show how to map accuracy on the predictable subset to full evaluations (Sec.
-4.4).
+In this section, we first introduce the problem formulation.
+Then we present the COD method in four parts, as illustrated in figure˜2: 1) We show the construction of sample-level difficulty scaling feature and present an improved mean-shift clustering algorithm (Sec.˜4.1); 2) We derive a performance scaling law corresponding to task difficulty variance, which supports extrapolating the performance-compute relationship for task clusters with similar difficulty features (Sec.˜4.2).
+We fit cluster-wise performance-compute curves on small models and filter extrapolatable clusters; 3) We extrapolate the performance on extrapolatable clusters and predict the accuracy of the target large model on the predictable subset (Sec.˜4.3); 4) We show how to map accuracy on the predictable subset to full evaluations (Sec.˜4.4).
+Problem Formulation.
+Consider a language model $M_{C}$ trained with a compute budget of $C$ measured in FLOPs.
+Let $\mathcal{P}$ be a set of downstream tasks that we aim to evaluate the model on.
+Each sample $T\in\mathcal{P}$ is defined by a question-answer pair $(q,a_{\text{true}})$.
+Given a question $q$, the model $M_{C}$ outputs a probability distribution $p(a|q;M_{C})$ over the space of all possible answers.
+Our goal is to predict the downstream task performance of a large language model $M_{C_{\text{target}}}$ using only evaluation results from smaller models $\{M_{C_{1}},M_{C_{2}},\ldots,M_{C_{n}}\}$ where $C_{i}\ll C_{\text{target}}$ for all $i$.
+Formally, we aim to found the prediction method $\phi$ to minimize the absolute prediction error over a group of tasks sets $\{\mathcal{P}_{j}\}_{m}$:
+$\arg\min_{\phi}\frac{1}{m}\sum_{i=1}^{m}\frac{1}{|\mathcal{P}_{j}|}\sum_{T\in\mathcal{P}_{j}}|\widehat{\text{Acc}}(C_{\text{target}},T)-\text{Acc}(C_{\text{target}})|,$
+$\widehat{\text{Acc}}(C_{\text{target}},T):=\phi\left(\{\text{Acc}(C_{i},T)\}_{i=1}^{n},\{C_{i}\}_{i=1}^{n},C_{\text{target}}\right),$
+where $\text{Acc}(C,T)$ denotes the accuracy of model $M_{C}$ on task $T$, and $\widehat{\text{Acc}}(C_{\text{target}},T)$ is the predicted accuracy for the target model.
 4.1 Clustering on Difficulty
@@
 We propose clustering tasks by similar performance scaling behaviors to minimize intra-cluster heterogeneity while ensuring robust evaluation with sufficient samples per cluster.
-To quantify task difficulty, we leverage the passrate(mean accuracy across multiple trials) from small models of progressively increasing scales.
-This process generates a feature vector for each task, which ideally exhibits monotonic growth, thereby reflecting the incremental enhancement of model capabilities.
-Improved clustering methods.
-We hope to adopt clustering algorithms with the following features: 1.
-Minimizing intra-class variance to ensure similar extrapolation properties within each cluster, 2.
-Automatic determination of cluster numbers, as the optimal number varies across evaluation sets and is difficult to pre-specify.
-While K-Means [27] requires pre-specifying cluster numbers and DBSCAN [11] may yield large intra-cluster variance, MeanShift [12] allows constraining this variance via a radius parameter, better suiting our needs.
+Specifically, we train a group of language models with increasing parameter counts.
+These models are trained with the same ratio of training tokens to compute per token.
+For each task, we generate 100 samples using top_p=0.7 and temperature=1.0 for each model, and compute the pass rate by averaging the results.
+This pass rate serves as an estimate of the model’s expected accuracy on the task.
+The resulting values are organized into a difficulty vector, ordered by increasing model size.
+For most tasks, this difficulty vector exhibits a monotonic increase, reflecting the gradual improvement of model capability with scale.
 Figure 3: t-SNE visualization of different clustering methods: DBSCAN(left), MeanShift(Middle), Improved-MeanShift(Right). Each point represents an evaluation sample.
-To further reduce intro-class variance, we propose an improved MeanShift algorithm to constrain the cluster diameter.
+After obtaining the difficulty feature vector for each task, we use the improved clustering algorithm that incorporates the following features: (1) Minimizing intra-class variance to ensure similar extrapolation properties within each cluster; (2) Automatic determination of cluster numbers, as the optimal number varies across evaluation sets and is difficult to pre-specify.
+To further reduce intra-class variance, we propose an improved MeanShift algorithm to constrain the cluster diameter.
 At the same time, we maintain a minimum number of tasks in each cluster to reduce metric fluctuations.
-We provide the t-SNE visualization of evaluation tasks on BBH [33].
+We provide the t-SNE visualization of clustering results evaluation tasks on BBH [38] to compare the proposed method and classic clustering algorithms including DBSCAN [14] and MeanShift [15].
 Each point represents an evaluation sample, and its color denotes the cluster type.
-As shown in Fig.
-3, our improved MeanShift effectively splits dense areas.
-We explain the details of clustering algorithms in Sec.
-A.1, and smoothing techniques in Sec.
-A.2.
+As shown in figure˜3, our improved MeanShift effectively splits dense areas whereas DBSCAN and the original MeanShift produce connected clusters with large within-cluster distances.
+We provide numerical comparison of clustering algorithms in Sec.˜5.3.1 and explain implementation details of imporved MeanShift in Appendix A.1, smoothing techniques in Appendix A.2.
 4.2 Fitting
 After clustering, we compute metrics for small models within each cluster.
-These small models are trained with the same ratio of training tokens to compute per token.
 We then apply a novel performance scaling law, based on theoretical analysis, to fit accuracy-to-compute curves for each cluster, focusing on clustered samples after excluding outliers.
@@
 Theorem 1 (Scaling Law for Downstream Task Performance).
-Given a language model trained with computational budget $C$, and a set of downstream tasks $P$, under the following assumptions:
-  1. 1.
-    Given a task $(q,\mathrm{ans})\in P$, the relationship between the answer loss and the compute follows a power law, $\frac{1}{n}\sum_{(q,\mathrm{ans})}\mathrm{loss}_{\mathrm{ans}}(C)\sim aC^{-b}+c$;
-  2. 2.
-    For tasks with a finite answer set, the model gives a random guess choice if it cannot truly solve it, i.e. $\mathbb{E}[\mathrm{Acc|incorrect}]=g$;
-  3. 3.
-    Task passrate equals the product of the predicted probability of each token $p_{\mathrm{ans}}=\prod_{t\in\mathrm{ans}}p(t)$, which means that each task has a unique answer and model output answer only without thinking progress.
-The expected accuracy on tasks $P$ can be modeled as:
-$\mathbb{E}_{p}[\mathrm{Acc}(C)]=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{\mu}\right)+o(\mu),$ (1)
-where $g$ represents the random guess performance floor, $1-g$ represents the maximum achievable performance improvement, $a,b,c$ are positive constants, and $\mu=\frac{1}{\#P}\sum_{(q,\mathrm{ans})\in P}\mathrm{loss}_{\mathrm{ans}}$; $\sigma^{2}=\frac{1}{\#P}\sum(\mathrm{loss}_{\mathrm{ans}_{t}}-\mu)^{2}$.
+Consider a language model $M_{C}$ trained with compute budget $C$ and a set of downstream tasks $\mathcal{P}$.
+Under the following assumptions: Assumption 1 (Power-law scaling of answer loss): the expected answer loss follows:
+$L_{\mathcal{P}}(C):=\mathbb{E}_{(q,a_{\text{true}})\sim\mathcal{P}}[L(q,a_{\text{true}};C)]=\alpha C^{-\beta}+\gamma$ (1)
+where $\alpha,\beta,\gamma>0$ are task-specific constants, with $\gamma$ representing the irreducible loss.
+Assumption 2 (Unique deterministic answers): Each question has a unique deterministic answer. The model receives score $1$ if and only if $M_{C}$ outputs $a_{\text{true}}$, and 0 otherwise.
+Assumption 3 (Accuracy decomposition): The expected accuracy decomposes as:
+$\mathbb{E}_{T\sim\mathcal{P}}[\mathrm{Acc}(C)]=g+(1-g)\cdot\mathbb{E}_{(q,a_{\text{true}})\sim\mathcal{P}}[p(a_{\text{true}}|q,M_{C})],$ (2)
+where $g\in[0,1]$ is the random guessing baseline.
+Then, the expected accuracy on task set $\mathcal{P}$ can be modeled as:
+$\mathbb{E}_{\mathcal{P}}[\mathrm{Acc}(C)]=g+(1-g)\left(\exp{(-\alpha C^{-\beta}-\gamma)}+\frac{\sigma_{L}^{2}(C)}{2\mu_{L}(C)}\right)+o\left(\sigma_{L}^{2}(C)\right)$ (3)
+where $\mu_{L}(C)=\mathbb{E}_{(q,a_{\text{true}})\sim\mathcal{P}}[L(q,a_{\text{true}};C)]$ is the mean loss and $\sigma_{L}^{2}(C)=\mathrm{Var}_{(q,a_{\text{true}})\sim\mathcal{P}}[L(q,a_{\text{true}};C)]$ is the loss variance across the task set.
 Proof Sketch.
-Assumption 3 (unique answers, no pre-answer reasoning) implies a negative logarithmic relationship between $\mathrm{loss}_{\mathrm{ans}}$ and individual task passrate, as correct task completion probability becomes the product of answer token probabilities.
-A key distinction, often overlooked, is how averaging occurs: the passrate metric averages $\exp(-\mathrm{loss}_{\mathrm{ans}})$ (an arithmetic mean), whereas applying the loss scaling law effectively averages loss before exponentiation (yielding a geometric mean).
-We show that the difference between these means can be approximated by $\frac{\sigma^{2}}{\mu}$ (where $\sigma^{2}$ and $\mu$ are the variance and mean of task passrates, respectively).
-Consequently, a performance scaling law derived from the loss scaling law is most valid for evaluation sets with limited difficulty variance.
-Our COD method achieves this by clustering tasks, thereby constraining intra-cluster difficulty variance and enabling better alignment with the performance scaling law.
-Finally, for tasks with a finite answer set, a random guessing score $g$ accounts for expected performance on unmastered tasks.
-Detailed proofs are provided in Appendix B.
+By definition of the language model loss, $p(a_{\text{true}}|q,M_{C})=\exp(-L(q,a_{\text{true}};C))$.
+Under Assumption 1, if the answer loss follows a power law $L\sim\alpha C^{-\beta}+\gamma$, then the task passrate should approximately scale as $\exp(-\alpha C^{-\beta}-\gamma)$.
+The key subtlety lies in the averaging: accuracy computes $\mathbb{E}[\exp(-L)]$ (arithmetic mean of passrates) while the loss scaling law gives us $\exp(-\mathbb{E}[L])$ (geometric mean).
+Using Taylor expansion:
+$\mathbb{E}[\exp(-L)]\approx\exp(-\mu_{L})\ +\frac{\sigma_{L}^{2}}{2\mu_{L}}$
+where $\mu_{L}$ and $\sigma_{L}^{2}$ are the mean and variance of the loss distribution.
+This approximation is accurate when tasks have similar difficulty feature ($\sigma_{L}^{2}/\mu_{L}^{2}\ll 1$), motivating our clustering approach to reduce intra-cluster variance.
+Assumption 3 adds the parameter $g$ for random guessing.
+The complete proofs are provided in Appendix B.
 ∎
-theorem 1 demonstrates that a metric of an evaluation set with similar difficulty features can be effectively modeled using the following formula:
+theorem˜1 demonstrates that a metric of an evaluation set with similar difficulty features can be effectively modeled using the following formula:
 $y(C)=g+(1-g)*e^{-aC^{-b}-c},$ (4)
 where $a$ and $b$ jointly influence how accuracy varies with $C$, $c$ controls the upper bound of the fitting curve, and $g$ represents the expected random guess metric for a task cluster.
-Parameters $a$, $b$, $c$, and $g$ are to be fitted.
-Note that these assumptions may not perfectly hold in practice, we provide additional discussions on assumption 3 in Appendix G.
+$a$, $b$, $c$, and $g$ are trainable parameters.
+Note that these assumptions may not perfectly hold in practice, we provide additional discussions on assumption 3 in Appendix I.
 4.3 Extrapolation
 To ensure reliable extrapolation, we identify clusters exhibiting robust scaling patterns, as some clusters may show saturated or non-emergent performance on smaller models, making them unsuitable for prediction.
+We aim to find an extrapolation subsets that represent the full set performance, and use the subset metric as a intermediate indicator for the prediction of the full set accuracy.
 A cluster is deemed extrapolatable if it meets two criteria: (1) its expected accuracy increases monotonically with model size, and (2) its performance converges to at least a predefined threshold $P$ (where $P\leq 1$ accounts for practical limits like ambiguous questions or finite training coverage).
-We filter out non-extrapolatable clusters using two rules based on the parameters from Eq. 2:
+We filter out non-extrapolatable clusters using two rules based on the parameters from Eq.˜4:
   1. 1.
@@
     Poor extrapolation reliability, indicated by an excessive $c$ value.
-In practice, for extrapolatable clusters, we set $a>1$, $b>0.1$, and $10\leq c<$ 1.
+In practice, for extrapolatable clusters, we set $a>1$, $b>0.1$, and $0\leq c<1$. Further ablation experiments are provided in Appendix C.1.
 The clusters satisfying these conditions form the predictable subset.
@@
 4.4 Mapping from Predictable Subset to Target Evaluation Set
-We map predictions from the predictable subset $T^{\prime}$ to the complete evaluation set $T$ using a function $f$.
+We map predictions from the predictable subset $\mathcal{P}^{\prime}$ to the complete evaluation set $\mathcal{P}$ using a function $g$.
 This approach is based on the observation that extrapolatable and non-extrapolatable samples, while differing in difficulty, often share question types, suggesting a preserved relative metric ordering.
-The mapping function $f:T^{\prime}\rightarrow T$ is continuous, smooth over $[0,1]$, monotonically increasing, and constrained to pass through $(0,0)$ and $(1,1)$.
-Empirical validation indicates a quartic function optimally captures this relationship:
-$f(x)=\alpha_{1}x^{4}+\alpha_{2}x^{3}+\alpha_{3}x^{2}+(1-\alpha_{1}-\alpha_{2}-\alpha_{3})x$ (3)
+The mapping function $f:\text{Acc}(\mathcal{P}^{\prime})\rightarrow\text{Acc}(\mathcal{P})$ is continuous, smooth over $[0,1]$, monotonically increasing, and constrained to pass through $(0,0)$ and $(1,1)$.
+Empirical validation indicates a polynomial function optimally captures this relationship:
+$f_{n}(x)=\alpha_{n-1}x^{n}+\alpha_{n-2}x^{n-1}+...+\alpha_{1}x^{2}+(1-\alpha_{1}-\alpha_{2}-...-\alpha_{n-1})x,$ (5)
+where $x$ represents the average accuracy of the predictable subset $\mathcal{P}^{\prime}$, and $n$ denotes the highest degree of the polynomial.
+In practical implementation, a quartic function is employed to fit the mapping (see Appendix C.2 for exhaustive experimental details).
 To ensure reliability, we calibrate $f$ using evaluation results from existing models as anchors.
-This subset-to-full mapping generally demonstrates robustness across diverse model architectures and training data, often permitting the use of external models (e.g., Qwen2-72B [40]) as anchors for many tasks (see Sec.
-5.3.3 for experiments).
-For data-sensitive tasks, however, models with similar training distributions provide more reliable anchors, prioritizing data consistency for mapping accuracy.
-The final metric prediction for a target LLM with estimated training computation $C_{0}$ is then $p=f\circ y(C_{0})$, combining the cluster-wise extrapolation $y(C_{0})$ from Eq.
-2 with the mapping $f$.
+This subset-to-full mapping generally demonstrates robustness across diverse model architectures and training data, often permitting the use of external models (e.g., Qwen2-72B [45]) as anchors for many tasks (see Appendix C.3 for experiments).
+However, for data-sensitive tasks, models with similar training distributions provide more reliable anchors, prioritizing data consistency for mapping accuracy.
+The final metric prediction for a target LLM with estimated training computation $C_{0}$ is then $p=f\circ y(C_{0})$, combining the cluster-wise extrapolation $y(C_{0})$ from Eq.˜4 with the mapping $f$.
 5 Experiments
 5.1 Experimental Setups
-In our experimental setup, we train several smaller versions of the target model architecture for prediction.
-These models vary in size, but share similar training procedures, with the training data scaled proportionally to their sizes.
-We show the training configurations of the model in Appendix C.
+In our experimental setup, we train nine language models ranging from 122M to 70B parameters in total, which share the same data distribution and architecture, with the training data scaled proportionally to their sizes.
+We show the detailed training configurations and recip in Appendix D.
 For evaluation, we adopt the following widely used benchmarks, including GSM8K [9], MATH [19], BBH [38], TriviaQA [24], MBPP [4], AGIEval [50], DROP [12], MMLU-pro [41].
-All models are evaluated in a few-shot in-context learning manner, and we aligned our evaluation setups with LLaMa3 [10].
-We evaluate our proposed COD performance scaling for LLMs against existing approaches on multiple public benchmarks.
-Using a series of smaller models with identical data distribution and architecture but varied configurations, we estimate the downstream task performance of a pretrained LLM with 70B parameters.
+All models are evaluated in a few-shot in-context learning manner, and we aligned our evaluation setups with LLaMa3 [13].We evaluate our proposed COD performance scaling for LLMs against existing approaches on multiple public benchmarks.
+Using eight smaller language models as known information, we estimate the downstream task performance of a pretrained LLM with 70B parameters.
 5.2 Prediction Experiments
-We compare COD against three representative prediction methods:
+We compare COD against four representative prediction methods:
   1. 1.
-    End-to-end performance-compute prediction: Directly extrapolates large model metrics from smaller model evaluation set metrics using performance scaling laws.
+    Loss-intermediate performance prediction [8]: First predicts the target LLM’s final training loss, then estimates downstream task metrics based on the relationship between smaller models’ evaluation metrics and their losses.
   2. 2.
-    Passrate-compute prediction [1, 20]: Estimates large model passrates from smaller model passrates.
+    End-to-end performance-compute prediction: Directly extrapolates large model metrics from smaller model evaluation set metrics using performance scaling laws [43].
+  3. 3.
+    Passrate-compute prediction [1, 23]: A variant of end-to-end method, which estimates large model passrates from smaller model passrates.
 We conduct 100 trials per evaluation set for smaller models to enhance reliability and report absolute prediction error on the passrate metric.
-  3. 3.
-    Loss-intermediate performance prediction [6]: First predicts the target LLM’s final training loss, then estimates downstream task metrics based on the relationship between smaller models’ evaluation metrics and their losses.
-Table 1: Absolute prediction error (in percentage points) on evaluation sets.
+  4. 4.
+    BNSL [7]: Another variant of end-to-end method, which decomposes the end-to-end mapping into a multi-segment power-law framework.
+Table 1: Absolute prediction error (%) on evaluation sets for predicting the performance of the 70B model.
 Errors < 2% are considered accurate (green), while errors > 5% are considered invalid (red).
@@
 […]
+Loss-intermediate 5.29 9.39 9.39 6.95 2.33 5.81 5.52 1.41 5.37 5.55
 End-to-end 3.10 6.00 4.00 3.86 0.64 0.68 1.75 6.00 4.11 3.72
 Passrate 5.02 8.80 6.71 8.80 3.51 4.00 7.34 6.78 0.26 2.74
-Loss-intermediate 5.29 9.39 9.39 6.95 2.33 5.81 5.52 1.41 5.37 5.55
+BNSL 5.17 13.05 4.23 5.88 13.05 5.86 2.55 0.82 1.53 7.42
 COD (w/o mapping) 2.24 5.26 4.70 0.50 2.91 1.98 0.89 5.26 1.08 0.57
@@
   1. 1.
-    COD (w/o mapping): Employs difficulty-based clustering, extrapolates within each cluster, and aggregates metrics, bypassing the subset-to-full mapping stage.
+    COD (w/o mapping): Employs difficulty-based clustering, extrapolates within all clusters, and aggregates metrics, bypassing the subset-to-full mapping stage.
   2. 2.
     COD (Complete): Our full proposed multi-stage approach, including clustering, predictable cluster filtering, subset extrapolation, and subset-to-full mapping.
-Comparative results are shown in Table 1.
+Comparative results are shown in Tab.˜1.
 Prediction accuracy is measured by the absolute error between predicted and actual performance.
@@
 While baselines perform well on some sets, their substantial errors on others undermine overall credibility.
-Figure 4: Performance-compute relationship for different prediction methods on BBH, MATH, and MMLU-pro evaluation sets.
-Fig.
-4 visualizes the performance-compute relationships.
+Figure 4: Performance-compute relationship for different prediction methods on eight evaluation sets.
+figure˜4 visualizes the performance-compute relationships.
+The COD method does not merely extend the existing scaling trend; instead, it effectively predicts whether growth slowdown will occur subsequently and enables better estimation of the magnitude of curve bending.
 On BBH, while all methods give similar final estimates, end-to-end and loss-intermediate methods fit small model data poorly.
@@
 On MATH and MMLU-pro, where predicting accelerated growth versus plateaus is crucial, the loss-intermediate method underestimates model ceilings, and the end-to-end method exceeds 3% error.
-COD’s superior performance stems from its nuanced analysis of difficulty distributions and scaling laws, allowing it to predict growth in challenging sets (e.g., MATH) and capture diminishing returns in saturated sets.
-The COD method is also capable of predicting LLMs’ performance after continual training phase on high-quality data.
-We conduct experiments and analysis on continual-trained LLMs and report prediction results in Appendix D.
+COD’s superior performance stems from its nuanced analysis of difficulty distributions and scaling laws, allowing it to predict growth in challenging sets and capture diminishing returns in saturated sets.
+To further validate the generalizability of the COD method, we examine its predictive capabilities across different model scales and for models after continual training.
+In addition to the 70B model, the COD method also demonstrates effectiveness in predicting the capabilities of the medium-size models, with experimental details and corresponding conclusions available in Appendix E.
+Beyond this, the method is further shown to be capable of predicting the performance of LLMs following a continual training phase on high-quality data.
+To validate this, we conducted dedicated experiments and in-depth analyses on continually trained LLMs, with results documented in Appendix F.
 5.3 Ablation Study
+In this section, we conduct ablation experiments on the clustering method and extrapolation function, while in the Appendix C we discuss ablation studies on the selection criteria for extrapolatable subsets, interpolation mapping methods, and the impact of anchor point selection on predictions.
 5.3.1 Comparison of Clustering Methods
-We assess the impact of different clustering algorithms on prediction accuracy.
-The goal is to achieve tight intra-cluster difficulty similarity (low average distance to center) while maintaining cluster stability (min.
-10 samples/cluster).
-We compare our Improved-MeanShift with DBScan, MeanShift, and K-Means.
-For K-Means, we adjust it to approximate our goals: (1) search for $k$ yielding min.
-cluster sizes $\approx 10$; (2) treat samples outside a radius (e.g., 2x average intra-cluster distance) from any cluster center as outliers, ensuring clusters don’t drop below 10 samples.
-We term this "Improved-KMeans" for this comparison.
-Clustering quality is measured by Intra-cluster Average Distance (IAD) and Outlier Rate (OR) (Table 2).
-Prediction benefits are measured by Extrapolation Error (EE) on the predictable subset and Final prediction Error (FE) (Table 3).
 […]
-Method MMLU-pro GSM8k MATH BBH
 […]
-K-Means 0.3236 - 0.2238 - 0.2238 - 0.6284 -
-DBScan 0.4242 0.56 0.5131 0.53 0.4775 0.68 0.7113 18.92
-MeanShift 0.2859 0.39 0.2852 0.61 0.2110 1.44 0.2679 20.72
-Improved-KMeans 0.1609 2.85 0.1321 2.73 0.0902 2.22 0.1953 37.23
-Improved-MeanShift 0.2225 4.40 0.1854 4.93 0.1463 2.66 0.2143 33.58
-[…]
-Method Mean Max MMLU-pro GSM8k MATH BBH
-[…]
-K-Means 3.62 3.76 8.16 8.99 3.69 3.69 0.01 0.00 2.62 2.34 8.16 8.99
-DBScan 3.93 4.08 4.38 4.36 3.72 3.69 4.08 4.12 4.38 4.16 3.53 4.36
-MeanShift 2.12 1.68 3.15 3.08 3.15 3.08 0.67 0.74 2.55 2.26 2.12 0.65
-Improved-KMeans 1.33 1.84 3.92 4.08 0.56 0.61 3.92 4.08 0.81 0.51 0.02 2.17
-Improved-MeanShift 1.23 1.66 2.20 2.23 1.27 1.35 2.20 2.23 1.14 1.28 0.29 1.77
-Table 2 shows Improved-KMeans and Improved-MeanShift yield better clustering (lower IAD) due to their intra-cluster distance constraints.
-Table 3 confirms these methods lead to lower EE and FE.
+Dataset K-Means DBScan MeanShift Improved-KMeans Improved-MeanShift
+IAD OR $E$ FE IAD OR $E$ FE IAD OR $E$ FE IAD OR $E$ FE IAD OR $E$ FE
+MMLU-pro 0.32 - 3.69 3.69 0.42 0.56 3.72 3.69 0.29 0.39 3.15 3.08 0.16 2.85 0.56 0.61 0.22 4.40 1.27 1.35
+GSM8k 0.22 - 0.01 0.00 0.51 0.53 4.08 4.12 0.29 0.61 0.67 0.74 0.13 2.73 3.92 4.08 0.19 4.93 2.20 2.23
+MATH 0.22 - 2.62 2.34 0.48 0.68 4.38 4.16 0.21 1.44 2.55 2.26 0.09 2.22 0.81 0.51 0.15 2.66 1.14 1.28
+BBH 0.63 - 8.16 8.99 0.71 18.92 3.53 4.36 0.27 20.72 2.12 0.65 0.20 37.23 0.02 2.17 0.21 33.58 0.29 1.77
+TriviaQA 0.44 - 2.97 2.46 0.70 6.38 1.11 0.81 0.25 6.77 3.64 4.90 0.12 11.97 1.18 1.12 0.19 11.54 1.58 1.64
+AGIEval 0.46 - 2.61 2.68 0.56 3.67 6.43 6.27 0.29 2.99 2.63 3.23 0.15 7.60 5.96 5.56 0.21 11.50 1.11 2.38
+DROP 0.56 - 1.66 1.64 0.67 11.08 3.03 2.66 0.25 11.81 4.18 4.00 0.14 21.42 3.99 5.24 0.20 19.88 0.26 0.23
+MBPP 0.34 - 2.53 2.67 0.51 12.80 1.57 1.41 0.22 15.60 2.40 1.22 0.17 19.40 2.39 3.25 0.17 21.60 1.56 2.19
+Mean - - 3.62 3.76 - - 3.93 4.08 - - 2.12 1.68 - - 1.33 1.84 - - 1.23 1.66
+Max - - 8.16 8.99 - - 4.38 4.36 - - 3.15 3.08 - - 3.92 4.08 - - 2.20 2.23
+Tab.˜2 shows Improved-KMeans and Improved-MeanShift yield better clustering (lower IAD) due to their intra-cluster distance constraints.
+The results also confirm these methods lead to lower EE and FE.
 Although Improved-KMeans has the best IAD, its GSM8k prediction is poor.
@@
 In contrast, our Improved-MeanShift, which automatically determines $k$ based on distance constraints, offers more stable clustering and the lowest maximum prediction error.
-Additional results in Sec.
-E.1 support these conclusions.
 5.3.2 Different Extrapolation Formulas
-Table 4: Ablation study of extrapolation formulas. EE, TR, FE shown for BBH, MATH, MMLU-pro.
+Table 3: Ablation study on extrapolation formulas. $E$, TR, FE shown for BBH, MATH, MMLU-pro.
 Method BBH MATH MMLU-pro
@@
 Ours ($f$) 0.29 52.46 1.77 1.14 81.24 1.28 1.27 94.38 1.35
-We ablate our proposed fitting formula $f(C)=g+(1-g)\cdot e^{-aC^{-b}-c}$ (Ours) by removing or modifying components: 1) $f_{1}(C)=e^{-aC^{-b}-c}$ (w/o random guess); 2) $f_{2}(C)=g+(1-g)\cdot e^{-aC^{-b}}$ (w/o constant $c$); 3) $f_{3}(C)=e^{-aC^{-b}}$ (Direct Power Law [20]).
-Table 4 shows Extrapolation Error (EE), Task Ratio of predictable subset (TR), and Final prediction Error (FE).
+We ablate our proposed fitting formula $f(C)=g+(1-g)\cdot e^{-aC^{-b}-c}$ (Ours) by removing or modifying components: (1) $f_{1}(C)=e^{-aC^{-b}-c}$ (w/o random guess); (2) $f_{2}(C)=g+(1-g)\cdot e^{-aC^{-b}}$ (w/o constant $c$); (3) $f_{3}(C)=e^{-aC^{-b}}$ (Direct Power Law [23]).
+Tab.˜3 shows Extrapolation Error (EE), Task Ratio of predictable subset (TR), and Final prediction Error (FE).
 Our proposed formula $f$ consistently achieves the lowest EE and FE.
@@
 The direct power law ($f_{3}$) fails to model the 0-1 metric range and the varying difficulty of improvement near random guess and saturation.
-TR’s minor influence on prediction error suggests robustness; even with low TR due to non-emergent tasks, their performance can be largely predicted via mapping (difficulty distributions visualized in Appendix F).
-5.3.3 Different Anchor Points in Interpolation Mapping
-We find that the mapping relationship from predictable subset metrics to full evaluation set metrics is similar across models with different training data and architectures.
-This allows leveraging pre-trained models as "anchors" to refine the mapping and improve final estimation accuracy.
-We use Qwen2-72B [40] (OOD anchor) and an in-house MoE model [23] with the same training data but different architecture (ID anchor).
-We first derive an interpolation curve using only small model metrics and fixed points $(0,0)$, $(1,1)$, then assess anchor compatibility.
-This shared mapping implies estimable subset metrics are highly correlated with full-set metrics and less prone to interference from other model parameters than loss-intermediate predictions.
-We test three configurations:
-  • COD (w/o anchor): Full COD pipeline, but no anchor points in the mapping phase;
-  • COD (w. OOD anchor): COD with Qwen2-72B as an out-of-distribution anchor;
-  • COD (w. ID anchor): COD with our in-house MoE model as an in-domain anchor.
-Table 5: Influence of anchor points in the mapping stage on prediction error (pp). Lower is better.
-Method Overall Individual Task Sets
-[…]
-w/o. anchor 3.96 8.80 2.17 5.46 5.08 1.68 8.80 2.38 4.44 1.68
-w. OOD anchor 1.59 2.56 2.22 0.94 1.84 1.04 2.56 1.86 1.18 1.04
-w. ID anchor 1.63 2.38 2.23 1.28 1.77 1.64 2.19 2.38 0.23 1.35
-Table 5 shows that incorporating either OOD or ID anchors consistently enhances prediction accuracy.
-This suggests a stable correlation between predictable subset and full-set metrics across diverse models, enabling the use of existing model evaluations to improve predictions for new models.
-Furthermore, since our clustering identifies intrinsic properties of evaluation sets, the derived predictable subsets are applicable to new models.
-Additional ablations on the interpolation method (Sec.
-E.2) confirms quartic functions are suitable.
+TR’s minor influence on prediction error suggests robustness; even with low TR due to non-emergent tasks, their performance can be largely predicted via mapping.
 6 Conclusion and Discussion
 In this work, we introduce a novel downstream performance scaling framework including (1) the COD method that effectively models the underlying diverse scaling patterns of evaluation tasks; (2) a scaling law for downstream task performance that provides a fitting formula for performance-compute extrapolation; and (3) a systematic methodology for identifying and leveraging predictable subset that provides a robust intermediate metric for accurate full-set performance predictions.
-We discuss the limitations of the COD method in Appendix G.
-Looking forward, our approach can be further expanded across model architectures, training methods, and evaluation set types, while extending this framework to address chain-of-thought reasoning patterns offer promising avenues for future research.
-Boarder Impact
-The LLMs performance scaling prediction method proposed in this work can be used to guide decision-making for training LLMs and to monitor whether the LLM training process aligns with expectations.
-Accurate prediction allows for better planning of compute and human resources, thereby playing a valuable role in scientifically advancing the training of larger LLMs.
-However, if the prediction results are inaccurate, it may lead to outcomes that do not meet expectations for the invested resources, potentially resulting in resource waste or economic losses.
+We discuss the limitations and future works in Appendix I.
+Ethics Statement
+We have read and adhered to the ICLR Code of Ethics.
+This work proposes a computational framework to enable more efficient resource allocation in the training of LLMs.
+The research is methodological in nature and aims to support more sustainable and responsible practices within the field of AI development.
+We provide a detailed account of our methods, theoretical proofs, and experimental settings.
+We openly discuss the limitations of our framework in Appendix I.
+This study does not involve human subjects or the use of sensitive personal data.
+We utilized language models at the writing level, including checking for grammatical errors in the article and modifying expressions.
+The use of language models had no impact on the article’s innovative contributions, experiments, or analytical perspectives.
+Reproducibility Statement
+We have made extensive efforts to ensure the reproducibility of our work.
+The core methodology, the Clustering-On-Difficulty (COD) framework, is detailed in Sec.˜4.
+The improved MeanShift clustering algorithm is described in Sec.˜4.1 , with full pseudocode provided in Appendix A.1 (Algorithm 1).
+Our performance scaling law (Theorem 1) is presented in Sec.˜4.2 , with a complete proof available in appendix˜B.
+We discuss the additional computational cost of the COD method in Appendix H.
+Our experimental setup, including model architectures, training data philosophy, and hyperparameters, is thoroughly documented in Sec.˜5.1 and Appendix D , with specific model configurations listed in Tab.˜A4.
+The evaluation benchmarks, protocols, and few-shot settings are described in Sec.˜5.1 and summarized in Tab.˜A5.
+Extensive ablation studies validating our component choices—including extrapolation formulas (Sec.˜5.3.2), clustering algorithms (Appendix 5.3.1), interpolation methods (Appendix C.2), and criteria for filtering clusters (Appendix C.1)—are provided to support our findings.
+We visualize the task difficulty distribution for each evalset in Appendix G.
 Appendix A Improvements of Clustering Algorithm
@@
 At the end of each iteration, we incorporate both the outliers from MeanShift and our marked unclustered samples into the next round of clustering, continuing this process until no further changes occur in sample labels.
-We present the pseudocode in algorithm 1.
+We present the pseudocode in algorithm˜1.
 Algorithm 1 Iterative MeanShift Clustering Algorithm
-1:Initialize all labels in the evaluation set to $-1$
+1:Initialize all labels in the evaluation set to 1-1
 2:repeat
-3:     Perform MeanShift clustering with radius $R$ on all samples labeled $-1$
+3:  Perform MeanShift clustering with radius $R$ on all samples labeled 1-1
 4:  Assign new labels to clustered samples
@@
 7:   if $\mathrm{dist}_{i}>R$ then
-8:              Reset label to $-1$
+8:     Reset label to 1-1
 9:   end if
@@
 11:  for each cluster do
-12:         if number of samples in cluster $<K$ then
-13:              Reset all samples in this cluster to $-1$
+12:   if number of samples in cluster <K<K then
+13:     Reset all samples in this cluster to 1-1
 14:   end if
@@
 17:until no label changes
+In the experiment, $K$ is empirically set to $10$, which has been verified through extensive experiments to be a reasonable and robust value.
+The fixed radius $R$ depends only on the value range and dimensionality (i.e., the number of predictive models) of the difficulty feature vectors.
+In our experiments, we fixed $R=0.3$ for all evaluation sets to ensure consistent overall clustering performance.
+For experiments with different numbers of small models used for fitting, the value of $R$ needs to be adaptively adjusted due to changes in the dimensionality of the feature vector.
+For instance, in the experiment of predicting the 12B-parameter model, $R$ was empirically reduced from $0.3$ to $0.2$ to ensure that the intra-class error is reasonable.
 Filtering zero-performance samples.
@@
 Appendix B Proof of Theorem 1
-We use lemma 1 to derive the scaling law for downstream task performance (theorem A1).
+We use lemma˜1 to derive the scaling law for downstream task performance (theorem˜1).
 Lemma 1 (Arithmetic-geometric mean difference).
@@
 Then the difference between the arithmetic mean and geometric mean can be estimated as:
-$\Delta$ $=\mu_{a}-\mu_{g}=\frac{1}{n}\sum_{i=1}^{n}x_{i}-\left(\prod_{i=1}^{n}x_{i}\right)^{\frac{1}{n}}=\frac{\sigma^{2}}{2\mu_{a}}+o(\mu_{a}).$ (4)
+$\Delta$ $=\mu_{a}-\mu_{g}=\frac{1}{n}\sum_{i=1}^{n}x_{i}-\left(\prod_{i=1}^{n}x_{i}\right)^{\frac{1}{n}}=\frac{\sigma^{2}}{2\mu_{a}}+o(\sigma^{2}).$ (6)
 Proof.
@@
 We can simplify:
-$\frac{1}{n}\sum_{i=1}^{n}\log x_{i}=$ $\log\mu+\frac{1}{n}\sum_{i=1}^{n}\left(\frac{(x_{i}-\mu_{a})}{\mu_{a}}-\frac{(x_{i}-\mu_{a})^{2}}{2\mu_{a}^{2}}\right)+o(\mu_{a})$
-$=$ $\log\mu+\frac{1}{\mu}\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}-\mu_{a}\right)}_{\text{equal to 0}}+\frac{1}{2\mu_{a}^{2}}\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu_{a})^{2}\right)}_{\sigma^{2}}+o(\mu_{a})$
-$=$ $\log\mu-\frac{\sigma^{2}}{2\mu^{2}}+o(\mu_{a}).$
+$\frac{1}{n}\sum_{i=1}^{n}\log x_{i}=$ $\log\mu+\frac{1}{n}\sum_{i=1}^{n}\left(\frac{(x_{i}-\mu_{a})}{\mu_{a}}-\frac{(x_{i}-\mu_{a})^{2}}{2\mu_{a}^{2}}+o\left((x_{i}-\mu_{a})^{2}\right)\right)$
+$=$ $\log\mu+\frac{1}{\mu}\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}-\mu_{a}\right)}_{\text{equal to 0}}+\frac{1}{2\mu_{a}^{2}}\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu_{a})^{2}\right)}_{\sigma^{2}}+o\left(\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu_{a})^{2}\right)$
+$=$ $\log\mu-\frac{\sigma^{2}}{2\mu^{2}}+o(\sigma^{2}).$
 Therefore:
-$\mu_{a}-\mu_{g}=\mu_{a}\left(1-\exp{\left(-\frac{\sigma^{2}}{2\mu^{2}}\right)}\right)+o(\mu_{a}).$ (7)
+$\mu_{a}-\mu_{g}=\mu_{a}\left(1-\exp{\left(-\frac{\sigma^{2}}{2\mu^{2}}\right)}\right)+o(\sigma^{2}).$ (9)
 When $\frac{\sigma^{2}}{2\mu^{2}}$ is small, this can be approximated as:
@@
 Theorem A1 (Scaling Law for Downstream Task Performance).
-Given a language model trained with computational budget $C$, and a set of downstream tasks $P$, under the following assumptions:
-  1. 1.
-    Given a task $(q,\mathrm{ans})\in P$, the relationship between the answer loss and the compute follows a power law, $\frac{1}{n}\sum_{(q,\mathrm{ans})}\mathrm{loss}_{\mathrm{ans}}(C)\sim aC^{-b}+c$;
-  2. 2.
-    For tasks with a finite answer set, the model gives a random guess choice if it cannot truly solve it, i.e. $\mathbb{E}[\mathrm{Acc|incorrect}]=g$;
-  3. 3.
-    Task passrate equals the product of the predicted probability of each token $p_{\mathrm{ans}}=\prod_{t\in\mathrm{ans}}p(t)$, which means that each task has a unique answer and model output answer only without thinking progress.
-The expected accuracy on tasks $P$ can be modeled as:
-$\mathbb{E}_{p}[\mathrm{Acc}(C)]=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{\mu}\right)+o(\mu),$ (9)
-where $g$ represents the random guess performance floor, $1-g$ represents the maximum achievable performance improvement, $a,b,c$ are positive constants, and $\mu=\frac{1}{\#P}\sum_{(q,\mathrm{ans})\in P}\mathrm{loss}_{\mathrm{ans}}$; $\sigma^{2}=\frac{1}{\#P}\sum(\mathrm{loss}_{\mathrm{ans}_{t}}-\mu)^{2}$.
+Consider a language model $M_{C}$ trained with compute budget $C$ and a set of downstream tasks $\mathcal{P}$.
+Under the following assumptions: Assumption 1 (Power-law scaling of answer loss): the expected answer loss follows:
+$L_{\mathcal{P}}(C):=\mathbb{E}_{(q,a_{\text{true}})\sim\mathcal{P}}[L(q,a_{\text{true}};C)]=\alpha C^{-\beta}+\gamma$ (11)
+where $\alpha,\beta,\gamma>0$ are task-specific constants, with $\gamma$ representing the irreducible loss.
+Assumption 2 (Unique deterministic answers): Each question has a unique deterministic answer. The model receives score $1$ if and only if $M_{C}$ outputs $a_{\text{true}}$, and 0 otherwise.
+Assumption 3 (Accuracy decomposition): The expected accuracy decomposes as:
+$\mathbb{E}_{T\sim\mathcal{P}}[\mathrm{Acc}(C)]=g+(1-g)\cdot\mathbb{E}_{(q,a_{\text{true}})\sim\mathcal{P}}[p(a_{\text{true}}|q,M_{C})],$ (12)
+where $g\in[0,1]$ is the random guessing baseline.
+Then, the expected accuracy on task set $\mathcal{P}$ can be modeled as:
+$\mathbb{E}_{\mathcal{P}}[\mathrm{Acc}(C)]=g+(1-g)\left(\exp{(-\alpha C^{-\beta}-\gamma)}+\frac{\sigma_{L}^{2}(C)}{2\mu_{L}(C)}\right)+o\left(\sigma_{L}^{2}(C)\right)$ (13)
+where $\mu_{L}(C)=\mathbb{E}_{(q,a_{\text{true}})\sim\mathcal{P}}[L(q,a_{\text{true}};C)]$ is the mean loss and $\sigma_{L}^{2}(C)=\mathrm{Var}_{(q,a_{\text{true}})\sim\mathcal{P}}[L(q,a_{\text{true}};C)]$ is the loss variance across the task set.
 Proof.
-We first use Assumption 3 to establish the relationship between model passrate and loss on a task:
-$-\log(p_{\mathrm{ans}})$ $=-\log\left(\prod_{t\in\mathrm{ans}}p(t)\right)=-\sum_{t\in\mathrm{ans}}\log(p_{t})=\mathrm{loss}_{\mathrm{ans}}.$ (10)
+For a task $T=(q,a_{\text{true}})\in\mathcal{P}$, under assumption 2, $a_{\text{true}}$ is deterministic and unique, thus $p(a_{\text{true}}|q,M_{C})$ can be decomposed into token-wise auto-regressive loss.
+$-\log(p(a_{\text{true}}|q,M_{C}))$ $=-\log\left(\prod_{i=1}^{k}p(t_{i}|q,t_{<i};M_{C})\right)$ (14)
+$=-\sum_{i=1}^{k}\log\left(p(t_{i}|q,t_{<i};M_{C})\right)$ (15)
+$=L(q,a_{\text{true}};C).$ (16)
 Then take the exponential of both sides, and then take the expectation with respect to different tasks in the evaluation set $p=(q,\mathrm{ans})\in P$.
 We note that both $p_{\mathrm{ans}}$ and $\mathrm{loss}_{\mathrm{ans}}$ are functions of $C$.
-$\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]$ $=\mathbb{E}_{p}[\exp(-\mathrm{loss}_{\mathrm{ans}}(C))]=\frac{1}{n}\sum_{(q,\mathrm{ans})\in P}\exp(-\mathrm{loss}_{\mathrm{ans}}(C)).$ (11)
-We can adopt lemma 1 to switch from arithmetic mean to geometric mean of $\mathrm{loss}$, and apply the power law assumption 1.
-$\frac{1}{n}\sum_{(q,\mathrm{ans}_{t})\in P}\exp(-\mathrm{loss}_{\mathrm{ans}}(C))=$ $\exp{\underbrace{\left(-\frac{1}{n}\sum_{(q,\mathrm{ans}_{t})\in P}\mathrm{loss}_{\mathrm{ans}}(C)\right)}_{\text{use loss scaling law}}}+\frac{\sigma^{2}}{2\mu}+o(\mu)$ (12)
-$=$ $\exp{(-aC^{-b}-c)}+\frac{\sigma^{2}}{2\mu}+o(\mu)$ (13)
-where $n=\#P$, and $\mu$, $\sigma^{2}$ follow definitions in the proposition.
-Finally, we use assumption 2 to align the answer passrate and the accuracy metric. We can adopt the law of total expectation:
-$\mathbb{E}_{p}[\mathrm{Acc}(C)]=$ $\mathbb{P}_{p}(\mathrm{correct})\mathbb{E}[\mathrm{Acc}|\mathrm{correct}]+\mathbb{P}_{p}(\mathrm{incorrect})\mathbb{E}[\mathrm{Acc|incorrect}]$ (14)
-Note that $\mathbb{P}_{p}(\mathrm{correct})=\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]$, $\mathbb{E}[\mathrm{Acc}|\mathrm{correct}]=1$ and $\mathbb{P}_{p}(\mathrm{incorrect})=1-\mathbb{P}_{p}(\mathrm{correct})$.
-We also define $g$ as the random guess accuracy performance, thus we have $\mathbb{E}[\mathrm{Acc|incorrect}]=g$.
-Take these results into Eq.
-14, and we have:
-$\mathbb{E}_{p}[\mathrm{Acc}(C)]$ $=\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]+(1-\mathbb{E}_{p}[p_{\mathrm{ans}}(C)])*g$ (15)
-$=g+(1-g)\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]$ (16)
-$=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{2\mu}\right)+o(\mu)$ (17)
-theorem A1 demonstrates that a metric of an evaluation set with similar difficulty features can be effectively modeled using the following formula:
-$f(C)=g+(1-g)\exp{(-aC^{-b}-c)}.$ (18)
-Appendix C Experiment Settings
-Model training.
-To establish performance predictions for large language models, we conduct systematic experiments with a suite of smaller-scale models across different parameter counts and training data volumes, while controlling for other training configurations such as learning rate, batch size, and additional hyperparameters.
-All models are trained on a constant learning rate scheduler with a few-step warmup stage and data with the same distribution.
-We list model configurations in Table A1.
+$\mathbb{E}_{p}[p(a_{\text{true}}|q,M_{C})]$ $=\mathbb{E}_{p}[\exp(-L(q,a_{\text{true}};C)]=\frac{1}{n}\sum_{(q,a_{\text{true}})\in P}\exp(-L(q,a_{\text{true}};C)).$ (17)
+We can adopt lemma˜1 to switch from arithmetic mean to geometric mean of $\mathrm{loss}$, and apply the power law assumption 1.
+$\frac{1}{n}\sum_{(q,a_{\text{true}})\in P}\exp(-L(q,a_{\text{true}};C))=$ $\exp{\underbrace{\left(-\frac{1}{n}\sum_{(q,a_{\text{true}})\in P}L(q,a_{\text{true}};C)\right)}_{\text{use loss scaling law}}}+\frac{\sigma_{L}^{2}(C)}{2\mu_{L}(C)}+o\left(\sigma_{L}^{2}(C)\right)$ (18)
+$=$ $\exp{(-\alpha C^{-\beta}-\gamma)}+\frac{\sigma_{L}^{2}(C)}{2\mu_{L}(C)}+o\left(\sigma_{L}^{2}(C)\right)$ (19)
+where $n$ in the number of tasks in $\mathcal{P}$, and $\mu$, $\sigma^{2}$ follow definitions in the proposition.
+Finally, we use assumption 3 to align the answer passrate and the accuracy metric.
+$\mathbb{E}_{T\sim\mathcal{P}}[\mathrm{Acc}(C)]$ $=g+(1-g)\cdot\mathbb{E}_{(q,a_{\text{true}})\sim\mathcal{P}}[p(a_{\text{true}}|q,M_{C})]$ (20)
+$=g+\frac{(1-g)}{n}\sum_{(q,a_{\text{true}})\in P}\exp(-L(q,a_{\text{true}};C))$ (21)
+$=g+(1-g)\left(\exp{(-\alpha C^{-\beta}-\gamma)}+\frac{\sigma_{L}^{2}(C)}{2\mu_{L}(C)}\right)+o\left(\sigma_{L}^{2}(C)\right)$ (22)
+Rationality of Assumption 3
+Assumption 3 is designed to accommodate tasks with finite answer sets.
+For such tasks, when calculating $\text{Acc}(C)$, possibilities outside the answer set are disregarded.
+When $p(a_{\text{true}}\mid q,M_{C})$ approaches 0, $\text{Acc}(C)$ is at the level of a random guess, $g$.
+When $p(a_{\text{true}}\mid q,M_{C})$ approaches 1, $\text{Acc}(C)$ is close to $p(a_{\text{true}}\mid q,M_{C})$.
+This assumption implies a linear relationship between $\text{Acc}(C)$ and the probability in the $(0,1)$ interval.
+The theorem itself is also effective for tasks with open answer sets, where the probability of a correct random guess can be assumed to be 0 (i.e., $g=0$).
+Appendix C Additional Ablation Studies
+C.1 The Criteria for Extrapolatable Subsets
+The criteria for fitting the extrapolation formula (Eq.˜4) are designed to ensure the following:
+  1. 1
+    $a>0$ and $b>0$: These ensure that accuracy is an increasing function of compute.
+Larger values of a and b signify that task performance scales more distinctly with compute, leading to fitting curves with better scaling properties and differentiability.
+  2. 2
+    $c>=0$: This ensures the extrapolated curve’s maximum value is less than or equal to $1$.
+An excessively large c implies that the fitting curve has a very low ceiling, which is characteristic of task subsets with poor scaling properties.
+These are thus considered non-extrapolatable.
+We conducted an ablation study on these parameters, as shown in Tab.˜A1.
+Starting from our baseline criteria ($a>1$, $b>0.1$, $0\leq c<1$), we individually relaxed the constraints on a, b, and c, and also observed the effect of removing the thresholds entirely.
+[…]
+Metric / Task Set Baseline ($a>1$, $b>0.1$, $0\leq c<1$) Ablate a ($a>0.5$) Ablate b ($b>0.05$) Ablate c ($0\leq c<0.5$) w/o threshold
+[…]
+Mean 1.18 1.63 1.18 1.67 1.19 1.63 1.44 1.97 2.39 2.36
+Max 2.20 2.23 2.20 2.23 2.20 2.23 2.88 2.99 6.24 5.97
+GSM8k 2.20 2.23 2.20 2.23 2.20 2.23 1.82 2.01 2.20 2.23
+MATH 1.14 1.28 1.14 1.28 1.14 1.28 1.14 1.28 0.80 0.79
+BBH 0.29 1.77 0.29 1.77 0.29 1.77 0.21 2.99 6.24 5.97
+TriviaQA 1.58 1.64 1.59 1.93 1.58 1.64 1.07 1.17 2.92 3.52
+MBPP 1.56 2.19 1.56 2.19 1.56 2.19 1.56 2.19 1.53 1.11
+AGIEval 1.11 2.38 1.11 2.38 1.11 2.38 0.08 1.59 2.22 2.65
+DROP 0.26 0.23 0.26 0.23 0.32 0.18 2.72 1.68 2.20 1.57
+MMLU-pro 1.27 1.35 1.27 1.35 1.27 1.35 2.88 2.86 1.00 1.03
+When the thresholds are removed entirely, the prediction performance degrades significantly.
+This is because numerous task clusters with poor scaling properties are included in the extrapolation, impairing the overall result.
+In contrast, individually relaxing the thresholds for a, b, or c still largely preserves the integrity of the filtering criteria.
+The performance shows only a slight drop compared to the baseline, indicating that while the filtering step is important, our method is not overly sensitive to the specific threshold values.
+C.2 Interpolation Method
+To evaluate different interpolation methods for prediction accuracy, we compared various mathematical approaches.
+Our baseline method uses quartic polynomial interpolation, which we compare against several alternative approaches, including Cubic spline interpolation, Cubic polynomial interpolation, and Quintic polynomial interpolation
+The comparative results across different benchmarks are shown in Tab.˜A2. We report the prediction error between the real performance of a large model and the mapping result.
+Table A2: Comparison of different interpolation methods across benchmarks.
+Prediction Error BBH Math MMLU-pro
+Cubic Spline 0.68 1.31 1.37
+Cubic Polynomial 3.38 1.12 1.35
+Quintic Polynomial 0.18 1.42 1.36
+Quartic Polynomial 1.77 1.28 1.35
+Furthermore, in figure˜A1, we plot the mapping process using different interpolation formulas, where the x-axis represents the predictable subset indices and the y-axis represents the full set indices.
+The red points are the numerical values to be fitted, green points represent predicted values, purple points represent anchor points, and blue points show the actual performance.
+The prediction performance shows certain robustness across different interpolation methods.
+We aim to use the simplest possible interpolation function while maintaining low prediction errors.
+Based on the above results, we observe that the Cubic Polynomial shows larger prediction errors due to underfitting on BBH, while the Cubic Spline exhibits some overfitting.
+Both Quartic Polynomial and Quintic Polynomial perform well, therefore, we chose the Quartic Polynomial method as it requires fewer fitting parameters.
+Figure A1: Performance mapping with different interpolation methods on the BBH evaluation set.
+The cubic spline is overfitted, and the cubic polynomial method is underfitted.
+Quartic polynomials and quintic polynomials are comparable, while a quartic polynomial has fewer parameters.
+When the number of small models used for fitting varies, the polynomial form also needs to be adaptively adjusted to ensure the robustness of the mapping.
+For instance, in the experiment of predicting the 12B model, the scaling function was adjusted from a quartic polynomial to a cubic polynomial to mitigate overfitting.
+C.3 Different Anchor Points in Interpolation Mapping
+We find that the mapping relationship from predictable subset metrics to full evaluation set metrics is similar across models with different training data and architectures.
+This allows leveraging pre-trained models as "anchors" to refine the mapping and improve final estimation accuracy.
+We use Qwen2-72B [45] (OOD anchor) and an in-house MoE model [26] with the same training data but different architecture (ID anchor).
+We first derive an interpolation curve using only small model metrics and fixed points $(0,0)$, $(1,1)$, then assess anchor compatibility.
+This shared mapping implies estimable subset metrics are highly correlated with full-set metrics and less prone to interference from other model parameters than loss-intermediate predictions.
+We test three configurations:
+  • COD (w/o anchor): Full COD pipeline, but no anchor points in the mapping phase;
+  • COD (w. OOD anchor): COD with Qwen2-72B as an out-of-distribution anchor;
+  • COD (w. ID anchor): COD with our in-house MoE model as an in-domain anchor.
+Table A3: Influence of anchor points in the mapping stage on prediction error ($p$). Lower is better.
+Method Overall Individual Task Sets
+[…]
+w/o. anchor 3.96 8.80 2.17 5.46 5.08 1.68 8.80 2.38 4.44 1.68
+w. OOD anchor 1.59 2.56 2.22 0.94 1.84 1.04 2.56 1.86 1.18 1.04
+w. ID anchor 1.63 2.38 2.23 1.28 1.77 1.64 2.19 2.38 0.23 1.35
+Tab.˜A3 shows that incorporating either OOD or ID anchors consistently enhances prediction accuracy.
+This suggests a stable correlation between predictable subset and full-set metrics across diverse models, enabling the use of existing model evaluations to improve predictions for new models.
+Furthermore, since our clustering identifies intrinsic properties of evaluation sets, the derived predictable subsets are applicable to new models.
+Additional ablations on the interpolation method (Sec.˜C.2) confirms quartic functions are suitable.
+Appendix D Experimental Settings and Training Recipe
+Training recipe.
+To establish performance predictions for large language models, we conduct systematic experiments with a suite of smaller-scale models across different parameter counts.
+All our models are trained from scratch on a corpus of text data.
+We do not fix the data budget for all models; instead, we maintain a consistent Data-to-CPT (Compute per token) ratio for all models, as mentioned in Sec.˜5.1.
+We list model configurations in Tab.˜A4.
+We adopt the in-house training data that comprises multilingual text corpora, with increased weighting for domains such as STEM, code, and general knowledge, following Llama3 [17], Deepseek-v2 [28], Fineweb-EDU [29], etc.
+We apply several de-duplication methods and data cleaning mechanisms to each data source to ensure high-quality tokens.
+The model architecture is consistent with Llama3.1 [17], incorporating Grouped-Query Attention (GQA) [2], SwiGLU activation function [35], RMSNorm [48] with Pre-normalization, etc.
+The models are trained using $BF16$ precision with a sequence length of $8192$ and a RoPE [37] base of $500,000$.
+We employ the AdamW optimizer with $\beta=(0.9,0.95)$, a weight decay of $0.1$, and a dropout rate of $0.1$.
+All models are trained on a constant learning rate scheduler with a few-step warmup stage.
+To determine the learning rate and batch size, we adopt the hyperparameter scaling laws from Liu et al.
+[28].
+Specifically, the optimal learning rate $\eta_{\mathrm{opt}}$, and the optimal batch size $B_{\mathrm{opt}}$ are defined as power laws of the compute, measured in FLOPs: $\eta_{opt}=a_{1}\cdot C^{-b_{1}}$ and $B_{opt}=a_{2}\cdot C^{b_{2}}$, where $a_{1}$, $b_{1}$, $a_{2}$, $b_{2}$ are parameters to be fitted.
+We perform a grid search on our small models to identify their optimal learning rates and batch sizes, and then extrapolate these findings to the bigger models.
 Table A4: Model architecture specifications across different sizes.
@@
 KV Heads 8 10 12 14 16 20 8 12 8
-Evaluation settings.
-We conducted performance scaling estimation experiments across 8 major LLM evaluation sets.
+Training Resources.
+The 7B dense model is trained on 923B tokens, consuming 52,800 H800 GPU-hours.
+The computational resources used for the other models can be estimated proportionally based on their respective compute requirements.
+Evaluation settings and protocol.
+We conducted performance scaling estimation experiments across eight major LLM evaluation sets.
 These evaluation sets span a diverse range of capabilities, including Math, Reasoning, Knowledge, Coding, Reading, and general abilities.
 All pretrained LLMs were evaluated using a few-shot methodology to obtain the performance metrics.
-Detailed information is provided in Table A2.
+Detailed information is provided in Tab.˜A5.
+Our evaluation methodology aligns with that used for the Llama3 [17] pre-trained models.
+We assess the models’ capabilities directly through few-shot text completion tasks without any instruction tuning or Supervised Fine-Tuning (SFT).
+This evaluation method is chosen because even a small amount of SFT data can significantly influence performance on downstream tasks, thereby not reflecting the inherent capabilities of the pre-trained model itself.
+Software Framework. All models are trained using the Megatron framework. The evaluation code is an in-house implementation designed to be consistent with the Llama3 [17] evaluation methodology.
 Table A5: Information of evaluation datasets.
@@
 MMLU-pro Comprehensive 12,032 5
-Appendix D Performance Prediction for Continue-Pretrained LLMs
+Appendix E Additional Prediction Results for a Medium-size Model
+[…]
+For this experiment, clustering relied exclusively on the passrate@100 data from the 122M to 1.9B source models to prevent data leakage from the target model.
+The results in Tab.˜A6 show that COD is more robust and less demanding on the size of source models compared to the baselines.
+For instance, the baseline methods perform poorly on the BBH benchmark, In contrast, COD accurately captures the trend of accuracy growth as compute increases and yields precise extrapolations.
+We will include a plot of this curve in the main text to visually demonstrate the clear advantage of COD.
+The prediction errors of all methods for the 12B model are significantly higher than those for the 70B model reported in Sec.˜5.2.
+Extrapolation accuracy improves when larger source models are employed for fitting, as a greater number of samples have “emerged”.
+If the largest source model is excessively small, numerous samples remain “unemerged”, resulting in substantial errors when extrapolating to a significantly larger model.
+Table A6: Absolute prediction error (%) on evaluation sets for predicting the performance of the 12B model.
+Errors < 2% are considered accurate (green), while errors > 5% are considered invalid (red).
+$\downarrow$ indicates lower is better.
+Task Sets/Metrics Ground Truth Loss-intermediate End-to-end BNSL COD
+Prediction Error Prediction Error Prediction Error Prediction Error
+Overall Metrics
+Mean - - 14.92 - 10.52 - 17.42 - 4.68
+Max - - 23.82 - 24.43 - 28.26 - 10.39
+Individual Task Sets
+GSM8k 70.47 60.86 9.61 63.66 6.81 80.99 10.52 60.08 10.39
+MATH 30.85 41.53 10.68 34.36 3.51 49.26 18.41 28.48 2.37
+BBH 60.76 36.94 23.82 36.33 24.43 36.10 24.66 57.751 3.05
+TriviaQA 58.84 68.42 9.57 59.88 1.04 87.10 28.26 57.92 0.93
+MBPP 47.80 28.53 19.27 36.51 11.29 27.95 19.85 44.50 3.30
+AGIEval 44.13 33.91 10.22 35.40 8.74 34.12 10.01 47.53 3.40
+DROP 54.35 35.05 19.30 37.19 17.16 36.73 17.62 63.50 9.14
+MMLU-pro 37.60 20.74 16.86 26.41 11.20 27.59 10.01 42.42 4.82
+Appendix F Performance Prediction for Continue-Pretrained LLMs
 Leading industry pre-trained LLMs (e.g., Deepseek-v3 [10], Llama3 [17], Qwen-2.5 [44]) adopt the Continual Pre-Training (CPT) strategy of concentrating high-quality data towards the end of the pre-training process.
@@
 Therefore, we supplement this by providing metric predictions for the high-quality CPT phase.
-The relationship between model parameter scale and the volume of CPT tokens is listed in Table A1.
+The relationship between model parameter scale and the volume of CPT tokens is listed in Tab.˜A4.
 We conduct the same COD pipeline for CPT models.
@@
 MATH 49.29 52.68 3.39
-Results listed in Table A3 show that the proposed COD method achieves an average prediction error of 2.33%.
+Results listed in Tab.˜A7 show that the proposed COD method achieves an average prediction error of 2.33%.
 We observe that MATH, GSM8k, and TriviaQA exhibit relatively large prediction errors. We hypothesize that there are two main categories of reasons for this inaccuracy:
@@
 This leads to potential significant fluctuations in the metrics after the CPT phase, thereby diminishing the accuracy of extrapolating to larger models.
-Appendix E Additional Ablation Studies
-E.1 Comparison of Clustering Methods on Extra Evaluation Sets.
-We provide additional clustering evaluation results across the remaining evaluation sets in Table A4 and Table A5, which maintain consistency with the conclusions presented in the main text.
-Table A4: Clustering performance on advanced task benchmarks (IAD: Intra-cluster Average Distance, OR: Outlier Rate)
-Method TriviaQA AGIEval DROP MBPP
-[…]
-K-Means 0.4388 - 0.4572 - 0.5554 - 0.3383 -
-DBScan 0.7039 6.38 0.5591 3.67 0.6651 11.08 0.5060 12.80
-MeanShift 0.2521 6.77 0.2886 2.99 0.2507 11.81 0.2167 15.60
-Improved-KMeans 0.1239 11.97 0.1536 7.60 0.1428 21.42 0.1667 19.40
-Improved-MeanShift 0.1871 11.54 0.2100 11.50 0.1974 19.88 0.1745 21.60
-Table A5: Prediction errors on advanced task benchmarks.
-Method Mean Max TriviaQA AGIEval DROP MBPP
-EE FE EE FE EE FE EE FE EE FE EE FE
-K-Means 2.44 3.36 2.97 8.99 2.97 2.46 2.61 2.68 1.66 1.64 2.53 2.67
-DBScan 3.04 2.79 6.43 4.36 1.11 0.81 6.43 6.27 3.03 2.66 1.57 1.41
-MeanShift 3.21 2.34 4.18 4.90 3.64 4.90 2.63 3.23 4.18 4.00 2.40 1.22
-Improved-KMeans 3.38 3.80 5.96 5.56 1.18 1.12 5.96 5.56 3.99 5.24 2.39 3.25
-Improved-MeanShift 1.13 1.61 1.58 2.38 1.58 1.64 1.11 2.38 0.26 0.23 1.56 2.19
-E.2 Interpolation Method
-To evaluate different interpolation methods for prediction accuracy, we compared various mathematical approaches.
-Our baseline method uses quartic polynomial interpolation, which we compare against several alternative approaches, including Cubic spline interpolation, Cubic polynomial interpolation, and Quintic polynomial interpolation
-The comparative results across different benchmarks are shown in Table A6. We report the prediction error between the real performance of a large model and the mapping result.
-Table A6: Comparison of different interpolation methods across benchmarks.
-Prediction Error BBH Math MMLU-pro
-Cubic Spline 0.68 1.31 1.37
-Cubic Polynomial 3.38 1.12 1.35
-Quintic Polynomial 0.18 1.42 1.36
-Quartic Polynomial 1.77 1.28 1.35
-Furthermore, in Fig.
-A1, we plot the mapping process using different interpolation formulas, where the x-axis represents the predictable subset indices and the y-axis represents the full set indices.
-The red points are the numerical values to be fitted, green points represent predicted values, purple points represent anchor points, and blue points show the actual performance.
-The prediction performance shows certain robustness across different interpolation methods.
-We aim to use the simplest possible interpolation function while maintaining low prediction errors.
-Based on the above results, we observe that the Cubic Polynomial shows larger prediction errors due to underfitting on BBH, while the Cubic Spline exhibits some overfitting.
-Both Quartic Polynomial and Quintic Polynomial perform well, therefore, we chose the Quartic Polynomial method as it requires fewer fitting parameters.
-Figure A1: Performance mapping with different interpolation methods on the BBH evaluation set.
-The cubic spline is overfitted, and the cubic polynomial method is underfitted.
-Quartic polynomials and quintic polynomials are comparable, while a quartic polynomial has fewer parameters.
-Appendix F Difficulty Distribution of Predictable Subset
+Appendix G Difficulty Distribution of Predictable Subset
 We analyze the proportion of predictable subset tasks across different difficulty levels.
-The difficulty distributions of predictable subset versus complete sets for different evaluation benchmarks are illustrated in Fig.
-A2.
+The difficulty distributions of predictable subset versus complete sets for different evaluation benchmarks are illustrated in figure˜A2.
 We use the scores from the 12B model as the basis for difficulty classification.
@@
 Figure A2: Difficulty distribution comparison on a 12B model between predictable subset and full evaluation set.
-Appendix G Limitations
+Appendix H Computational Cost
+The extra computational overhead of running COD is not expensive compared to training a series of small models with increasing parameter sizes.
+The main additional cost comes from performing 100 inference evaluations on the evaluation set for each small model (the cost of clustering algorithm is negligible compared to the inference cost).
+The computational complexity is $O(TMN)$, where $T$ is the number of evaluation runs, $M$ is the number of tokens for one evaluation, and $N$ is the maximum parameter size of the small models used for prediction.
+The corresponding token usage is $O(TM)$.
+In particular, for an evaluation set requiring $1M$ tokens, a total of $100M$ tokens for small model inference is needed.
+In our experiments, the training token count for the 12B small model is $1.554T$.
+Considering that a training token is typically 3 times more costly than an inference token, the additional cost of COD is approximately $\frac{100M}{1.554T*3}\approx 0.002\%$ of the training cost.
+Appendix I Limitations
 Influence of model structure.
@@
 Chain-of-thought performance prediction.
-theorem A1 assumes that evaluation sets directly assess models’ ability to provide answers.
+theorem˜1 assumes that evaluation sets directly assess models’ ability to provide answers.
 However, increasingly more evaluations allow models to think before providing answers.